### v1

# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

Abstract
This paper examines the impact of racial discrimination in hiring on employment, wage, and wealth disparities between black and white workers. Using a labor search-and-matching model with racially prejudiced and non-prejudiced firms, we show that labor market frictions sustain discriminatory practices as an equilibrium outcome. These practices account for 44% to 52% of the average wage gap and 16% of the median wealth gap. Discriminatory hiring also amplifies unemployment and wage volatility for black workers, increasing their labor market risks over business cycles. Eliminating prejudiced firms reduces these disparities and improves black workers’ welfare, though it slightly decreases overall economic welfare.
JEL classification: D14, E21, J15, J64, J65
Keywords: search-and-matching, heterogeneous agents, racial inequality, unemployment, wealth distribution, business cycles
1 Introduction
Discrimination is often considered inefficient or a response to information asymmetry. [^1] See Becker (1957) and Lang and Spitzer (2020) for discussions on taste-based and statistical discrimination theories. Market competition and better information are expected to eliminate discrimination over time. However, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities. [^2] For example, Couch and Fairlie (2010); Biddle and Hamermesh (2013); Kuhn et al. (2020); Derenoncourt and Montialoux (2021); Lippens et al. (2023). Resume studies consistently reveal racial discrimination even when controlling for candidate qualifications (e.g., Bertrand and Mullainathan, 2004; Bertrand and Duflo, 2017; Kline et al., 2022). Gaps in labor income and wealth between black and white households have persisted long after the Civil Rights Movement (e.g., Cajner et al., 2017; Derenoncourt et al., 2023). Small and Pager (2020) argues that discrimination reinforces itself across domains and constitutes a form of market failure. Only recently has macroeconomic research begun to explore racial disparities. [^3] See, for example, Nakajima (2021), Aliprantis et al. (2023), Boerma and Karabarbounis (2021), Ganong et al. (2020), Lee et al. (2021). Yet, the role of discriminatory practices in shaping these disparities remains underexplored. Developing a general equilibrium framework to explain the persistence of discrimination and its consequences is crucial for understanding the macroeconomic impacts of racial disparities and informing effective policymaking.
This paper develops a labor search-and-matching framework to capture the persistence of hiring discrimination based on race. We show that hiring discrimination accounts for 44% to 52% of the black-to-white average wage gap and 16% of the racial median wealth gap. Additionally, hiring discrimination intensifies unemployment and wage volatility for black workers, creating a countercyclical racial unemployment gap and diverging racial wage cyclical patterns. The model incorporates two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally. By construction, racial differences in the model are reflected as differences in bargaining power and the likelihood of experiencing wealth destruction. While all workers compete for jobs, white workers have additional opportunities with prejudiced firms. Matched worker-firm pairs negotiate wages to maximize their joint surplus. Upon calibration, the model endogenously produces lower job-finding rates, higher unemployment rates, and lower bargained wages for black workers. Even without additional financial frictions, the model generates significant racial disparities in wealth accumulation.
The main message of this paper is that labor market friction can sustain hiring discrimination as an equilibrium outcome. Our calibration shows that prejudiced firms incur higher vacancy posting costs to discriminate against black workers. As we gradually increase these costs, prejudiced firms slowly reduce their vacancy postings and employment, while their profits adjust to satisfy the free-entry condition. Ultimately, we more than double the benchmark calibrated posting cost until the entire expected value of entering as a discriminatory firm is eliminated. This indicates that prejudiced firms maintain sufficient profit to absorb higher vacancy posting costs, even under competitive market conditions.
Moreover, hiring discrimination perpetuates the equilibrium economy through firms’ vacancy postings. Prejudiced firms compete with non-prejudiced firms for unemployed white workers, driving up wage rates for white workers across the economy. This wage pressure reduces the expected profits of non-prejudiced firms, leading them to post fewer vacancies overall. The resulting decline in non-prejudiced job openings directly limits employment opportunities for black workers, contributing to lower wages, higher unemployment rates, and adverse wealth accumulation for them.
In a counterfactual economy without prejudiced firms, the racial unemployment gap disappears, and the wage gap decreases significantly. Black workers’ employment opportunities and wages increase because non-prejudiced firms expand their vacancy postings to fill positions left by prejudiced firms. As a result, black workers experience a welfare increase, with the largest gain among the high-productivity black workers. White workers, however, lose the exclusive job opportunities that prejudiced firms previously provided. Their average wage declines and their unemployment rate rises. This leads to a moderate welfare decrease, especially among the lowest-productivity and lowest-wealth white workers. Due to asymmetrical job destruction shocks and vacancy posting costs, non-prejudiced firms cannot fully replace prejudiced firms. However, if we maintain the race-specific job destruction shocks, eliminating prejudiced hiring has a much smaller adverse impact on white workers.
We further introduce aggregate fluctuations to the economy and show that hiring discrimination leads to distinct business cycle dynamics for black and white workers. Our model successfully replicates the observed countercyclical and volatile racial unemployment rate gap, as well as the more procyclical and volatile average wage rate for black workers. Upon removing prejudiced firms, the business cycle patterns of the racial unemployment rate gap disappear. The volatility and cyclicality of average wages for black and white workers become much more similar. This provides theoretical evidence to Cajner et al. (2017) that firm-side discrimination contributes to the business cycle differences between black and white workers, and non-market demographic factors alone do not fully account.
Lastly, we examine the business cycle welfare change between black and white workers when discriminatory hiring contributes to more volatile labor market outcomes for black workers. We compare the average consumption equivalence of workers during business cycles after removing prejudiced firms. Not all workers benefit from a less volatile economy without discriminatory firms. Black workers gain an average of 0.02% in welfare, whereas white workers face a welfare loss of 0.01%. Different from the steady-state welfare values, among black workers, those with high productivity and high wealth gain the most. Among white workers, those with high productivity and high wealth lose the most, experiencing the largest welfare decreases within their group.
This paper contributes to the ongoing discussion on the persistence of racial inequality by focusing on the disparate labor market conditions of white and black workers. Numerous studies have documented racial discrimination in pay and employment opportunities that persist despite policy interventions (e.g., Coate and Loury, 1993; Black, 1995; Rosén, 1997; Bertrand and Mullainathan, 2004; Manduca, 2018). Recent empirical research stresses that racial wage and employment gaps persist across education levels, skill sets, and cohorts. These disparities are not solely attributable to differences in individual qualifications (e.g., Pena, 2018; Cheng et al., 2019). Building on this insight, our paper develops a model in which discriminatory hiring practices endogenously perpetuate labor market inequalities. Our work adds to the literature by providing a theoretical framework that explains how hiring discrimination can be sustained as an equilibrium outcome in the labor market.
While much of the literature focuses on static disparities, fewer studies document how macroeconomic fluctuations impact black and white workers differently. Notably, Couch and Fairlie (2010) demonstrate that black workers are disproportionately affected during economic downturns, being the last hired in upturns and the first fired in downturns. Biddle and Hamermesh (2013) find that the discriminatory wage gap is procyclical, widening during economic expansions. Similarly, Cajner et al. (2017) show that black workers experience higher unemployment rate volatility and are more likely to be involuntarily employed part-time. Daly et al. (2020) suggest that limited employment opportunities for black workers contribute to the growing racial earnings gap. Our paper adds to this body of work by evaluating the impact of hiring discrimination on the cyclical aspects of racial disparities, examining how discriminatory practices exacerbate volatility and cyclicality in employment and wages for black workers.
An emerging strand of literature explores racial disparities in wealth accumulation. Studies such as Derenoncourt et al. (2023), Derenoncourt et al. (2022), Kuhn et al. (2020), Barsky et al. (2002), and McIntosh et al. (2020) document significant wealth gaps between black and white households. For example, Derenoncourt et al. (2023) provide a historical account of wealth segregation over the past 150 years, showing persistent and substantial disparities. Boerma and Karabarbounis (2021) and Aliprantis et al. (2023) analyze the impact of historical discrimination on earnings, bequests, and capital returns within steady-state models that do not consider aggregate risks. Furthermore, Ganong et al. (2020) demonstrate that income risks affect individuals differently across racial groups due to wealth differences. Building on these findings, Bartscher et al. (2021) and Lee et al. (2021) discuss how monetary policy can have disparate effects on workers of different racial backgrounds. Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over into long-term wealth accumulation disparities between black and white workers.
Germane to our project, Nakajima (2021) develop a search-and-matching model to examine how monetary policies exacerbate racial differences in the labor market. Our model explicitly distinguishes between discriminatory and non-discriminatory firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy. We analyze the spillover effects of discriminatory hiring practices on both labor and wealth disparities, offering a comprehensive theoretical understanding of these dynamics over the long run and across business cycles.
In a broader context, our paper joins the growing discussion on the distributional impact of economic growth and macroeconomic policies by focusing on the heterogeneous outcomes for black and white workers. This aligns with studies like Caballero and Hammour (1994), Jaimovich and Siu (2020), and Heathcote et al. (2020), which provide evidence that recessions disproportionately hurt disadvantaged individuals. Borella et al. (2018) show that incorporating gender differences in a life-cycle model improves its empirical fit, emphasizing the importance of accounting for demographic heterogeneity. Krusell and Smith (1998) conclude that heterogeneity in wealth does not alter business cycle fluctuations. Yet, Yum et al. (2023) show that heterogeneity could generate large aggregate fluctuations when introducing non-convexity in budget constraint through progressive tax. Our paper demonstrates that incorporating racial differences into labor market models enhances their ability to capture business cycle fluctuations. It also provides insights into how racial wedges in labor search processes transmit individual risks and heterogeneity into aggregate economic dynamics.
The rest of the paper proceeds as follows. Section 2 lays out theoretical framework. Section 3 discusses the calibration strategy. Section 4 examines the steady-state implications of racial discrimination. Section 5 provides business cycle implications. Section 6 concludes the paper.
2 Model
In this section, we present a model of labor market discrimination without aggregate uncertainty. We use this framework to understand the role of hiring discrimination in driving the racial gaps in labor market outcomes and wealth.
2.1 Environment
The model has a unit measure of workers, who are either black or white, $R\in\{bl,wh\}$. Two types of firms post vacancies to hire unemployed workers, subject to search and matching frictions. Prejudiced firms ($p$) actively discriminate against black workers and hire only white workers. Non-prejudiced firms ($np$) hire both black and white workers without discrimination. Prejudiced and non-prejudiced firms also have type-specific job destruction shocks. To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote the variables in the next period.
Workers are either employed (with $p$ or $np$ firm) or unemployed. Those who become unemployed in the current period receive unemployment benefits and continue to receive them in the future with probability $P_{e}$. The workers face idiosyncratic productivity shocks $s$, following an AR(1) process $\log(s^{\prime})=\rho_{s}\log(s)+\epsilon_{s}$, with $\epsilon_{s}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{s}^{2})$. Workers also differ in terms of their asset holdings. They have access to risk-free assets, and the workers can save using these assets to partially insure themselves against the labor market risks. Following Mukoyama (2013), workers also receive race-specific extreme wealth shocks with probability $\epsilon_{R}\in\{\epsilon_{bl},\epsilon_{wh}\}$. Upon realization of the shock, the worker loses all their wealth. Taken together, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$), and the endogenous distribution of workers is given by $\mu(e,R,s,a)$. All endogenous aggregate variables depend on $\mu$. Individual workers may move around the distribution between periods; however, the overall distribution remains stationary. We omit the record of $\mu$ in the steady state and will reintroduce $\mu$ when augmenting the model with aggregate shock.
2.2 Labor market search and matching
The total number of unemployed workers $u$ is the sum of unemployed black ($u_{bl}$) and white ($u_{wh}$) workers. The number of $np$ vacancies available is $v_{np}$, while the number of $p$ firm vacancies is $v_{p}$. Since the non-prejudiced firms search among both black and white unemployed workers, the non-prejudiced market tightness is given by $\theta_{np}=v_{np}/u$. On the other hand, prejudiced firms employ only white unemployed workers, and hence, the prejudiced market tightness is $\theta_{p}=v_{p}/u_{wh}$. Following Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018), unemployed workers and vacant firms match via a constant return to scale matching function
$M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}},$ (1)
with $\iota>0$. As documented by Den Haan et al. (2000), this functional form ensures that the matching probabilities lie within 0 and 1. The probability for an unemployed worker to match with a vacant $np$ firm is $f(\theta_{np})=M(u,v_{np})/u=(1+\theta_{np}^{-\iota})^{-1/\iota}$, while the probability that a white unemployed worker matches with a vacant $p$ firm is $f(\theta_{p})=M(u_{wh},v_{p})/u_{wh}=(1+\theta_{p}^{-\iota})^{-1/\iota}$. Correspondingly, the probability of filling a vacant $np$ firm is $q(\theta_{np})=M(u,v_{np})/v_{np}=(1+\theta_{np}^{\iota})^{-1/\iota}$, while the probability of filling a vacant $p$ firm is $q(\theta_{p})=M(u_{wh},v_{p})/v_{p}=(1+\theta_{p}^{\iota})^{-1/\iota}$. In addition, non-prejudiced matches get separated with probability $\lambda_{np}$, while the prejudiced matches separate at the rate $\lambda_{p}$.
2.3 Unemployment Insurance
Unemployment insurance is characterized by the replacement rate $h$, probability of eligibility $P_{e}$, and maximum eligibility level $\chi$. Following Setty and Yedid-Levi (2021), eligible workers receive unemployment benefits $b(R,s,a)=min\{h\bar{w}(R,s,a),\chi\}$, where $\bar{w}(R,s,a)$ is the counterfactual wage earned by an employed worker with race $R$, productivity $s$, and wealth $a$. We adopt the counterfactual wage to ease the computation burden of tracking wage history. Similar to Mitman and Rabinovich (2015), newly unemployed workers receive unemployment benefits with certainty and continue to receive benefits next period with probability $P_{e}$. If an unemployed worker loses their eligibility to receive benefits, they continue to remain ineligible in the future as well. Unemployment benefits are funded through proportional tax $\tau$ on the labor income, and the government sets $\tau$ to balance its budget. [^4] We intentionally model a more realistic and complex unemployment insurance structure to capture the racial disparities in the incidence and take-up of unemployment insurance. This helps us generate racial differences in income and wealth distribution, given the asymmetric labor market and wealth risks.
2.4 Workers
The value function of an employed individual with race $R$, productivity $s$, asset $a$, and working with a $np$ firm is given by $W_{np}(R,s,a)$, while that of a white worker employed with a $p$ firm is given by $W_{p}(wh,s,a)$. Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced ($U^{I}_{np}(R,s,a)$) or a prejudiced ($U^{I}_{p}(wh,s,a)$) firm previously. An unemployed worker who is not eligible for unemployment benefits earns $U^{N}(R,s,a)$ over their lifetime. All the workers discount their future utility by $\beta$. Similar to Nakajima (2012) and Setty and Yedid-Levi (2021), we assume that workers cannot borrow. This imposes an exogenous constraint of $a^{\prime}\geq 0$ on all workers.
2.4.1 Employed with $np$ firm
Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in an $np$ firm. The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their lifetime discounted utility. Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$. The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of extreme wealth shock $\epsilon_{R}$.
$W_{np}(R,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})\hat{W}_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}\hat{W}_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (2)
s.t.
$c+a^{\prime}$ $=(1-\tau)\omega_{np}(R,s,a)+(1+r)a+d$
The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3.
$\hat{W}_{np}(R,s^{\prime},a^{\prime})=\lambda_{np}U^{I}_{np}(R,s^{\prime},a^{\prime})+(1-\lambda_{np})W_{np}(R,s^{\prime},a^{\prime})$ (3)
If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$. Otherwise, one may stay employed with probability $1-\lambda_{np}$. If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
2.4.2 Employed with $p$ firm
Prejudiced firms $p$ hire only white workers. The rest of the problem is similar to the previous case and is given by Equation 4.
$W_{p}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{W}_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{W}_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (4)
s.t.
$c+a^{\prime}$ $=(1-\tau)\omega_{p}(wh,s,a)+(1+r)a+d$
The matched worker receives a job destruction shock specific to $p$ firms, with probability $\lambda_{p}$. A worker losing the job in the current period is eligible for unemployment benefits and earns value $U^{I}_{p}$ in the next period. For each $s^{\prime}$ and after realizing $\epsilon_{wh}$, an employed worker’s future value expands to Equation 5.
$\hat{W}_{p}(wh,s^{\prime},a^{\prime})=\lambda_{p}U^{I}_{p}(wh,s^{\prime},a^{\prime})+(1-\lambda_{p})W_{p}(wh,s^{\prime},a^{\prime})$ (5)
2.4.3 Unemployed and eligible workers
Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or an $np$ firm. In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
Black worker from a $np$ firm
Equation 6 describes the value of an unemployed black worker eligible for benefits. The worker receives unemployment insurance payout, $b_{np}(bl,s,a)$, net of labor income tax rate $\tau$, previous savings and lump-sum dividend transfer $d$. The worker encounters the extreme wealth shock with probability $\epsilon_{bl}$.
$U^{I}_{np}(bl,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{bl})\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})+\epsilon_{bl}\hat{U^{I}}_{np}(bl,s^{\prime},0)\big{]}\Big{\}}$ (6)
s.t.
$c+a^{\prime}$ $=(1-\tau)b_{np}(bl,s,a)+(1+r)a+d$
Their future value $\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})$ expands according to Equation 7:
$\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))[P_{e}U^{I}_{np}(bl,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(bl,s^{\prime},a^{\prime})]$ (7)
The unemployed worker finds a job with probability $f(\theta_{np})$. If the worker doesn’t get matched, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $1-P_{e}$.
White worker from a $np$ firm
Unemployed white workers are recruited by both $p$ and $np$ firms. Their value function is described by Equation 8. Like a black worker, a white worker has cash-on-hand from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer. They also encounter an extreme wealth shock with probability $\epsilon_{wh}$.
$U^{I}_{np}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U^{I}}_{np}(wh,s^{\prime},0)\big{]}\Big{\}}$ (8)
s.t.
$c+a^{\prime}$ $=(1-\tau)b_{np}(wh,s,a)+(1+r)a+d.$
Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$. They can receive a job offer from a $p$ firm with probability $(1-f(\theta_{np}))f(\theta_{p})$, an offer from an $np$ firm with probability $f(\theta_{np})(1-f(\theta_{p}))$, and receive both offers together with probability $f(\theta_{np})f(\theta_{p})$. The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$) and continue to receive unemployment benefits with probability $P_{e}$. If the worker receives both $p$ and $np$ offers, they will choose the job with higher lifetime utility.
$\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (9)
$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{np}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
White worker from a $p$ firm
Equation 10 describes the problem faced by a white worker who last worked with a $p$ firm. The problem is identical to the previous case, except that they receive a payout of $b_{p}(wh,s,a)$ from the unemployment insurance.
$U^{I}_{p}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U^{I}}_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (10)
s.t.
$c+a^{\prime}$ $=(1-\tau)b_{p}(wh,s,a)+(1+r)a+d$
After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the worker’s future value is described in Equation 11. It is nearly identical to Equation 9 except for the value of remaining unemployed from a $p$ firm rather than a $np$ firm.
$\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (11)
$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
2.4.4 Unemployed and ineligible workers
Ineligible workers have lost their unemployment insurance and are still searching for jobs. They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again. Thus, their income comes only from their previous savings and lump-sum dividends and does not depend on their prior employment situation.
Black worker
Equation 12 describes the problem of a black worker searching for jobs without unemployment insurance.
$U^{N}(bl,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{bl})\hat{U^{N}}(bl,s^{\prime},a^{\prime})+\epsilon_{bl}\hat{U^{N}}(bl,s^{\prime},0)\Big{]}\Big{\}}$ (12)
s.t.
$c+a^{\prime}$ $=(1+r)a+d$
And Equation 13 expands the future value of the black worker.
$\hat{U^{N}}(bl,s^{\prime},a^{\prime})$ $=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))U^{N}(bl,s^{\prime},a^{\prime})$ (13)
White worker
Equation 14 describes the problem of a white worker searching for jobs without unemployment insurance.
$U^{N}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{wh})\hat{U^{N}}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U^{N}}(wh,s^{\prime},0)\Big{]}\Big{\}}$ (14)
s.t.
$c+a^{\prime}$ $=(1+r)a+d$
And Equation 15 expands the future value of the white worker in this scenario.
$\hat{U^{N}}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (15)
$+\underbrace{f(\theta_{p})f(\theta_{np})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth np and p sector jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing unemployed}}$
Thus, the problem faced by an ineligible worker is analogous to that faced by an eligible worker, except that they don’t receive unemployment benefits.
2.5 Firms
There are a large number of $p$ and $np$ firms that post vacancies after paying the vacancy posting costs. The firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
2.5.1 Vacant $np$ firm
Vacant $np$ firms pay a cost of $\kappa_{np}$ and search among all unemployed workers irrespective of their race. These firms match with an unemployed worker with probability $q(\theta_{np})$. Equation 16 describes the value of a vacant $np$ firm, $V_{np}$.
$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}}_{\text{value of matching with atype (s,a) black worker}}\Big{]}$ (16)
$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{np provides higher value than a p firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}}_{\text{value of matching with (s,a) white worker}}\Big{)}$
$+\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{p provides higher value than a np firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))}_{\text{value of matching (s,a) worker without competing p offer}}\Big{]}\Big{\}}da,$
The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$. A $np$ firm can match with either a black or a white worker from the current unemployment pool. $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers. Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ are the densities of the unemployed black and white workers over $s$ and $a$. These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$. Since the white unemployed workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ match and begin producing only if it gives them a higher value or if the worker did not receive a $p$ firm offer. The last two rows of Equation 16 reflect the cases that a white worker will produce with the $np$ firm.
2.5.2 Vacant $p$ firm
Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers. The probability of matching with a white unemployed worker is $q(\theta_{p})$. Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}}_{\text{value of matching with (s,a) white worker, $p$ offers higher value than a np firm}}$ (17)
$+\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))}_{\text{value of matching with (s,a) white worker without competing np offer}}\Big{]}da$
Since the vacant $p$ firm searches only among white workers, the probability of matching with an unemployed white worker having productivity $s$ and asset $a$ is given by $\phi_{u}(wh,s,a)/u_{wh}$. Like the case of $np$ firms, the white worker accepts the $p$ match only if the $p$ offer is more favorable or if they don’t receive a competing $np$ offer.
We assume that there is free entry of firms, and hence both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
2.5.3 Producing $np$ firm
The value of a producing $np$ firm is described by Equation 18. $j(R,s,a)$ describes the period profit. The worker’s productivity $s$ enters the production function. The matched firm produces $sf(k)$ units of output, pays the rental and the depreciation cost of capital, $(r+\delta)k$, and the wage cost, $\omega_{np}(R,s,a)$.
$J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\Big{(}\frac{1-\lambda^{np}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
where
$j(R,s,a)$ $=sf(k)-(r+\delta)k-\omega_{np}(R,s,a).$
The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$. Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm. The future value of the firm incorporates this probability.
2.6 Producing $p$ firm
Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm. The firm hires capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
$J_{p}(wh,s,a)$ $=\max_{k}\Big{\{}j(wh,s,a)+\Big{(}\frac{1-\lambda^{p}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})J_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}J_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (19)
where
$j(wh,s,a)$ $=sf(k)-(r+\delta)k-\omega_{p}(wh,s,a).$
2.7 Wage bargaining
In each period, matched worker-firm pairs bargain for wages over the matched surplus. A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$. The worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$. The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
$\omega_{np}(R,s,a)$ $=\operatorname*{arg\,max}_{\omega_{np}}\Big{(}W_{np}(R,s,a)-U^{I}_{np}(R,s,a)\Big{)}^{\xi_{R}}J_{np}(Ra,s,a)^{1-\xi_{R}}$ (20)
$\omega_{p}(wh,s,a)$ $=\operatorname*{arg\,max}_{\omega_{p}}(W_{p}(wh,s,a)-U^{I}_{p}(wh,s,a))^{\xi_{wh}}J_{p}(wh,s,a)^{1-\xi_{wh}}$ (21)
We define the stationary equilibrium in Appendix A.
3 Calibration
A period in the model represents a quarter. We calibrate the benchmark model to match the relevant U.S. economy moments. We have two sets of parameters. The first group of parameters is chosen externally based on literature and empirical evidence without using model-generated data, while the second group of parameters is calibrated internally by simulating our model to match a set of relevant data moments. We set the share of black workers to be 19.1% in the model, with the remaining being white workers. Table 1 shows the internally calibrated parameter values and their targeted moments, as well as the externally chosen parameters and their sources.
Table 1: Calibration and targeted statistics
Parameter Value Description Target statistics data model
Chosen internally
$\beta$ 0.9943 discount factor K/Y 10.26 10.26
$\iota$ 1.3012 matching elasticity job finding rate - black 0.4946 0.4946
$\kappa_{p}$ 4.2622 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1 1
$\lambda_{np}$ 0.0644 np sector job destruction shock job separation rate - black 0.0644 0.0644
$\lambda_{p}$ 0.0268 p sector job destruction shock job separation rate - white 0.0380 0.0380
$\xi_{bl}$ 0.1397 bargaining power - black mean wage ratio 0.75 0.75
$\xi_{wh}$ 0.2110 bargaining power - white firm profit share 0.033 0.033
$\epsilon_{bl}$ 0.0179 extreme wealth shock - black zero wealth share - black 0.18 0.18
$\epsilon_{wh}$ 0.0086 extreme wealth shock - white zero wealth share - white 0.07 0.07
Chosen externally
$\alpha$ 0.6600 elasticity of labor matching Nakajima (2012)
$\theta_{n}$ 0.2890 capital share of output Nakajima (2012)
$\delta$ 0.0150 quarterly depreciation rate Nakajima (2012)
$\rho_{s}$ 0.9411 persistence of idiosyncratic shock PSID
$\sigma_{s}$ 0.1680 standard deviation of idiosyncratic shock PSID
$h$ 0.4000 UI replacement rate Mitman and Rabinovich (2015)
$\xi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
$Pe$ 0.5385 probability of UI eligibility maximum weeks of eligibility
Notes: This table reports the parameters, their values, and descriptions. The top panel presents the parameters chosen internally by minimizing the distance between model generated moments and data. The last two columns of the top panel compare the targeted moments in the data and in the model simulations. The bottom panel reports the parameters chosen externally, their values, and description.
3.1 Preferences
We set the period utility function $u(c)$ to be $\log(c)$. The discount factor, $\beta$, is calibrated to match the quarterly capital-output ratio of 10.26, the value used by a number of studies including Den Haan et al. (2010) and Carroll et al. (2017). The resulting value of $\beta$ is 0.9943, and the corresponding quarterly real interest rate is 1.3%.
3.2 Production
The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha}$. We choose $\alpha$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015 following Nakajima (2012).
3.3 Productivity and wealth shocks
We use hourly real wage from the Panel Study of Income Dynamics (PSID) to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process. Our estimation strategy closely follows Setty and Yedid-Levi (2021). We run a standard Mincer wage regression of log wage on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of males aged 25 years and above. We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data. The estimated regression is at a biennial frequency, given that the PSID data is available once in two years. We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process. The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency. We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.
Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth. The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers. They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth. (Nakajima, 2021).
3.4 Unemployment insurance
The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$. We choose the replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others. Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model. The eligibility probability $P_{e}$ is chosen to be 0.5385 to generate an average unemployment benefit duration of 26 weeks, as in Mitman and Rabinovich (2015).
3.5 Labor search and wage bargain
The labor market turnover statistics are obtained from Cajner et al. (2017). The elasticity of the matching function, $\iota$, targets the job finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018). We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers respectively. The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively. [^5] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers. On the other hand, white workers work in both $p$ and $np$ firms, and hence $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380). The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, to be 1 following Wolcott (2021). [^6] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$. We find that the prejudiced firms pay a lot more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated to be 4.2622, compared to $\kappa_{np}$ taking a value of 2.735. The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021). The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012). Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.211.
3.6 Assessing the model as a quantitative theory of racial disparity
Our calibration successfully captures the racial gaps along three important dimensions: income, wealth, and labor market outcomes. We now discuss each of these dimensions in more detail before proceeding with the quantitative exercises examining the macroeconomic impact of racial discrimination.
(a) Labor income share
(b) Wealth share
Figure 1: Labor Income and Wealth Distribution
Note: This figure compares the steady state model generated distributions of wealth and labor income with their empirical analogs. The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Survey of Consumer Finance (SCF), respectively.
First, the model successfully captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we don’t target these distributions explicitly. As with the majority of incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile. On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
Table 2: Steady State Racial Inequality
Moments Data Model
Unemp rate (Black) 0.12 0.12
Unemp rate (White) 0.05 0.05
Mean wealth ratio 0.23 0.28
Median wealth ratio 0.17 0.33
Note: This table compares the steady state wealth and unemployment moments between black and white workers with the corresponding empirical data.
Second, the model is able to capture the disparities in unemployment, labor income, and wealth between black and white workers. As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth. In addition to the targeted moments, Table 2 demonstrates that our model successfully captures the untargeted employment and wealth moments. Specifically, our model reproduces the 5% unemployment rate among white workers compared to 12% among black workers. Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to the prejudiced sector. Our model results show that the $p$ sector with the equilibrium market tightness of 0.4, though smaller than the $np$ market tightness of 0.73, provides an important advantage for white workers by increasing their job-finding rate compared to black workers. On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al. (2014). However, the model understates the median wealth gap between black and white workers. The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in the data (Kaplan et al., 2014).
In sum, the model captures the racial inequalities across labor income, wealth, and unemployment outcomes. We next explore how racial discrimination in hiring contributes to these differences.
4 Steady state results
In this section, we first establish how a search framework sustains hiring discrimination in equilibrium. Then, we examine the steady-state impact of racial discrimination by comparing the benchmark economy to an alternative economy without hiring discrimination. Lastly, we explore the heterogeneous welfare implications associated with eliminating hiring discrimination.
4.1 Sustaining racial discrimination
The labor search-and-matching framework allows firms to earn profit from the surplus generated by the search frictions. This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium. It contrasts with a frictionless canonical neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957). In particular, our benchmark model has $p$ firms retaining a profit of 0.04 while $np$ firms retain a profit of 0.03 in the steady state (see Column 1 of Table 3).
Next, we examine the steady-state interplay between discriminatory and non-discriminatory firms. Since discriminatory hiring remains profitable in the benchmark equilibrium, we gradually increase penalties by raising the vacancy posting costs for discriminatory firms until they are no longer profitable. Figure 2 displays the dynamics of employment, vacancy, and profits of the prejudiced and non-prejudiced firms. As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment. Conversely, with fewer $p$ firms competing in the market, $np$ firms see a steady increase in profits, vacancy postings, and employment.
It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks. As $p$ firms’ profit declines, the profit of $np$ firms increases at a slower rate. Additionally, as $np$ firms replace $p$ firms, they recruit from both black and white workers, rather than only white workers displaced from $p$ firms. This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
In the limit, as $\kappa_{p}$ approaches the total expected value of the $p$ sector, discriminatory ($p$) firms almost completely exit the market, since posting vacancies exclusively for white workers becomes unprofitable. The value of $\kappa_{p}$ at which $p$ firms exit is 217% of the benchmark value of 4.26. Meanwhile, non-discriminatory ($np$) firms become nearly the sole employers in the economy. The gray dotted lines in Figure 2 illustrate a counterfactual steady state where we entirely remove $p$ firms from the benchmark model. Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of completely eliminating $p$ firms.
Figure 2: Penalizing discrimination
Note: This figure displays the dynamics between $p$ and $np$ firms as $\kappa_{p}$ increases. The top panel plots the change in profit, the middle panel plots the change in vacancy posting, and the bottom panel plots the change in employment. The red dotted line depicts the dynamics of $p$ firms, the blue solid line depicts the dynamics of $np$ firms, and the gray dotted line depicts the limit where no $p$ firm exists. The horizontal axes record the values of $\kappa_{p}$ from the benchmark $\kappa_{p}$ at 4.2622 to 217% of the benchmark value, at 9.2444. The vertical axes record the model-generated value of each variable.
4.2 Impact of hiring discrimination on racial inequality
In this subsection, we compare the benchmark economy to one without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring. Table 3 presents the comparisons. All firms in the counterfactual models provide equal hiring opportunities to black and white workers. Column 2 removes the $p$ firms from the model, which leaves all workers with equal job destruction shocks at the benchmark $\lambda_{np}$. In Column 3, we re-calibrate the job destruction shocks to race-specific and have $\lambda_{wh}$ and $\lambda_{bl}$ as race-specific job destruction shocks. [^7] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates. This counterfactual economy only has one type of firm. We assign white workers a job destruction shock, $\lambda_{wh}$, to be 0.038 and black workers a destruction shock, $\lambda_{bl}$, to be 0.0644 to keep black and white workers’ job separation rates as in the calibration. Re-assigning the job destruction shocks further isolates the impact of job separation.
Table 3: Impact of hiring discrimination
Moments Benchmark Eliminate $p$ firms Race specific $\lambda$
Households
job separation rate - black 0.06 0.06 0.06
job separation rate - white 0.04 0.06 0.04
job finding rate - black 0.50 0.59 0.62
job finding rate - white 0.66 0.59 0.62
unemp rate - black 0.12 0.10 0.09
unemp rate - white 0.05 0.10 0.06
mean wage - black 1.60 1.96 1.82
mean wage - white 2.13 2.22 2.11
mean wage ratio 0.75 0.88 0.86
mean wealth ratio 0.28 0.33 0.33
median wealth ratio 0.33 0.44 0.44
Firms
$p$ firm profit 0.04 - -
np firm profit 0.03 0.06 0.06
p firm vacancy 0.02 - -
np firm vacancy 0.05 0.10 0.07
p firm employment 0.54 - -
np firm employment 0.40 0.90 0.94
Labor Market
$p$ market tightness 0.40 - -
$np$ market tightness 0.73 1.02 1.12
Aggregate Outcomes
Y 3.06 2.95 3.06
K/Y 10.26 10.22 10.25
average wage 2.04 1.96 2.05
unemp rate 0.07 0.10 0.06
Note: This table compares the benchmark steady state with a counterfactual steady state without $p$ firms. In the counterfactual model, all firms provide equal hiring opportunities to black and white workers and keep race-specific job destruction shock. All other parameters stay the same as benchmark calibration.
The first panel shows the changes in racial employment, wage, and wealth disparities after removing prejudiced hiring. The first two rows describe the job separation rates. By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the difference as the benchmark. The next two rows describe the job-finding rates. Since there is no $p$ sector in the economy in Columns 2 and 3, the job-finding rates are equalized for black and white workers. White workers only have one sector for employment. Hence, their job-finding rate reduces. Black workers have more job opportunities as the $np$ sector posts more vacancies; hence, their job-finding rate increases. Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2. Altogether, it reduces the racial gap in the unemployment rate. In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes. In Column 3, the unemployment rate gap closes from 7% to 3% because of the remaining racial difference in $\lambda$.
Turning to the wage impact, black workers can bargain for a higher wage rate as they receive more job opportunities when we remove discriminatory hiring. White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rate. As a result, the average black-to-white wage ratio increases from 75% in benchmark to 88% in Column 2 and to 86% in Column 3. We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44. The result remains stable between Columns 2 and 3. We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
The remaining gaps in wage and wealth come from the bargaining power and extreme wealth shock disparities between black and white workers. [^8] Though not the primary purpose of this paper, we present the results after removing racial differences in bargaining power and extreme wealth shock in Appendix B.
The second panel displays the changes in firm attributes. We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock. Without discriminatory firms, $np$ firms’ profit doubles, vacancy posting, and employment level increase. These changes result in the aforementioned advantages for black workers in the counterfactual economies. In Column 3, white workers’ lower job destruction shock drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
The third panel of Table 3 displays the labor market conditions. In response to the $p$ firms exiting, more $np$ firms endogenously enter the market. The labor market tightness of the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms and 1.12 when we split $\lambda$ by race. The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
The last panel of Table 3 describes the aggregate outcomes. With more $np$ firms entering, the aggregate output stays relatively unchanged in Column 3. The capital-output ratio slightly decreases to 10.25. The economy-wide average wage increases to 2.05 from benchmark 2.04, and the aggregate unemployment rate decreases to 6%. However, Column 2 experiences a more significant reduction in output, average wage, and higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs rather than a resource for production.
Our results show that the presence of discriminatory firms can explain sizable fractions of income and wealth inequality. Penalizing prejudiced firms primarily improves the employment opportunities and wage outcomes for the black workers; it also worsens the situation for the white workers. Removing $p$ firms reduces the job-finding rate and, hence, the unemployment rate for white workers. In addition, they also lower the value of their outside option in wage bargaining, which, in turn, affects their income and wealth outcomes. To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by the black and the white workers.
4.3 Welfare analysis
We follow Krusell et al. (2010) in measuring the welfare change for black and white workers by calculating the change in average consumption equivalence after removing discriminatory hiring. [^9] Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$. Under an alternate economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state. We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$. With $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$. We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change. We aggregate over the counterfactual distribution rather than benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms. Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift. We present the average welfare change between black and white workers by productivity types and by wealth quintiles in Table 4.
Overall, the economy experiences a 5.8% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race. Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%. This is because of the rise in their job opportunities and higher wage rates. On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%. This is because of the decrease in their wage rate and the increase in the unemployment rate.
Mid-panels in Table 4 present the heterogeneous welfare change for black and white workers by productivity types. Though the welfare reduction of eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms. In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race. In comparison, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
Table 4: Heterogeneous welfare change
Average welfare gain (%) Eliminate $p$ firms Race specific $\lambda$
Overall -5.83 1.17
Black White Black White
6.89 -8.83 10.94 -1.14
by productivity
Low 6.30 -9.03 10.31 -1.07
Mid 6.92 -8.83 10.98 -1.14
High 7.41 -8.61 11.49 -1.19
by wealth
Low 20% 6.85 -9.43 10.98 -1.13
40-60% 6.95 -8.95 10.94 -1.17
Top 20% 6.88 -8.14 10.72 -1.09
Note: This table compares the average consumption equivalence change from the benchmark steady state to the equilibrium where we eliminate $p$ firms (Columns 1 and 2) and further adjust the job destruction shock to be race-specific (Columns 3 and 4). Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The bottom panels of Table 4 present the welfare change by wealth quintiles for black and white workers. Removing discriminatory firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race. The most considerable reduction in welfare (9.43%) happens for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
5 Business cycle dynamics
In this section, we first document the racial disparities over the business cycle. Then, we introduce aggregate uncertainty in our benchmark model to quantify the role of hiring discrimination in generating disparate dynamics in the business cycle. Lastly, we discuss the welfare implications of hiring discrimination over the business cycle.
5.1 Racial disparities over the business cycles
Table 5 summarizes the business cycle disparities in the labor market between black and white workers. The top panel presents the dynamics of the racial unemployment rate gap, and the bottom panel presents the dynamics of the average wage. Cyclicality is measured by the correlation between the variable and real GDP. Volatility is calculated by the standard deviations of the variable relative to the standard deviation of real GDP.
Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference. We then log the values and apply the HP filter with a smoothing parameter 1600 to extract cyclical components. We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions. It has a relative volatility of 9.71.
Table 5: Business cycle statistics, US data, 1996-2014
Cyclicality Volatility
Unemployment rate
Black-White gap -0.6255 9.7100
Average wage
Black 2.6041 2.5352
White 1.2729 2.0848
Note: This table provides business cycle statistics for the US. The unemployment rate was obtained from the Current Population Survey (CPS) from 1996 to 2014. Average wage statistics are constructed from the Panel Study of Income Dynamics (PSID) of the same period. All variables are logged and HP filtered with a smoothing parameter of 1600.
Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage. This is because of the composition change of workers over the business cycles. Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994). For the same reason, we report the cyclical properties of the average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index. The wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics. The wage volatility is calculated by the relative standard deviations of the year-to-year change in wage from the wage regression. [^10] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors. Specifically, we follow Bils (1985), Solon et al. (1994), and Devereux (2001) and run the regression: $\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it}$ (22) where $Y_{t}$ refers to real GDP. Demographic variables $x_{it}$ include gender, education, labor market experience, and marital status. We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$. We measure cyclicality using coefficient $\beta$ from the regression. The volatility is the standard deviation of the year fixed effects, $\nu_{t}$. We then adjust the estimation to quarterly frequency. We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics. Black workers’ average wage is more volatile (2.54) than white workers (2.08).
5.2 Augmented model with aggregate shocks
We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2. The TFP shocks follows an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$. The output of the matched firm-worker pairs depends on the realizations of both aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers across employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$). The next period’s aggregate distribution $\mu^{\prime}$ is determined by $(z,\mu)$, and the law of motion is given by $\mu^{\prime}=\Gamma(z,\mu)$.
We follow Cooley et al. (1995) and Boppart et al. (2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015. All other parameters stay the same as in Table 1. This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
We obtain the stochastic equilibrium of our model by using the sequence space method of Boppart et al. (2018). We first solve non-linearly for the perfect foresight transitions to a single small MIT shock, i.e., an unexpected shock to the aggregate TFP. We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock. Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycle as a linear combination of the impulse response and the TFP realizations. [^11] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system. In Appendix C, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
Table 6: Business cycle model vs data
Cyclicality Volatility
Data Model Data Model
Unemployment rate
Black-White gap -0.6255 -0.6731 9.7100 0.3009
Average wage
Black 2.6041 1.0749 2.5352 1.0755
White 1.2729 0.9984 2.0848 0.9985
Note: This table compares the data moments and the benchmark generated moments on the black-white unemployment and wage gaps over the business cycles.
Table 6 compares data and model moments on unemployment and wages. Similar to standard search models, the stochastic dynamics of our model are under the influence of the Shimer Puzzle (Shimer, 2005). The augmented benchmark model generates similar cyclicality in the unemployment rate gap as data but much less volatility in the unemployment rate gap. Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility. However, qualitatively, we have a higher wage cyclicality and volatility for black workers than for white workers.
5.3 Impact of hiring discrimination over business cycle
We compare the benchmark dynamics with the counterfactual model without p firms to examine the role of discriminatory hiring over aggregate shocks. [^12] As with Table 3, the counterfactual economy is a one-sector economy with only $np$ firms. We do not further adjust race-specific destruction shocks.
5.3.1 Impulse responses
Figure 3 plots the response of black and white workers’ labor market, wealth, and consumption outcomes after a 1% unanticipated expansionary TFP shock. We compare the benchmark model response to the counterfactual model without $p$ firms. Overall, our models generate countercyclical unemployment rates and procyclical wage, wealth, and consumption movements.
Figure 3: Impulse responses
Note: This figure plots the impulse response functions of unemployment rates, black-white average wage and wealth ratios, and aggregate consumption between black and white workers to a one percent increase in aggregate TFP at date zero. The red solid line denotes the responses from the benchmark model, and the blue dashed line denotes the responses from the counterfactual model without $p$ firms.
In the model without $p$ firms, black workers’ unemployment rate decreases less than that of the benchmark economy after the expansionary shock. This implies that their unemployment rate is less volatile without $p$ firms. This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities to all workers. For white workers, the difference is much smaller between the benchmark and the model without $p$ firms. Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model. This is because the $p$ sector has a low job destruction shock. Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without discriminatory firms. This suggests more volatile procyclical wage and wealth responses for black workers than for white workers with the presence of $p$ firms. Lastly, We don’t observe a discerning difference in the consumption responses with and without $p$ firms.
5.3.2 Stochastic simulation
We simulate the stochastic equilibrium of the economy over 10,000 periods under aggregate TFP shocks, comparing outcomes with and without $p$ firms. While impulse responses to an MIT shock capture the transition path following a one-time disturbance, the stochastic equilibrium focuses on the economy’s short-term responses to continuous aggregate shocks. The results are summarized in Table 7.
The counterfactual equilibrium without $p$ firms eliminates the racial unemployment rate gap, resulting in the disappearance of its cyclical properties. Similarly, the racial differences in the cyclicality and volatility of average wages are largely reduced. As shown in Table 7, in the benchmark model, the black-white unemployment rate gap exhibits strong countercyclicality (-0.6731) and volatility (0.3009), both of which vanish in the absence of $p$ firms. For average wages, the cyclicality for black workers decreases from 1.0749 in the benchmark to 1.0131 without $p$ firms, while for white workers it increases slightly from 0.9984 to 1.0115. Similarly, the volatility for black workers falls from 1.0755 to 1.0139, while for white workers it increases from 0.9985 to 1.0116. This convergence demonstrates that hiring discrimination is a primary driver of the labor market’s cyclical disparities between black and white workers. These findings align with Cajner et al. (2017), who argue that demographic differences alone cannot fully account for observed racial disparities over the business cycle.
Table 7: Impact of hiring discrimination on business cycle racial gaps
Cyclicality Volatility
Benchmark No p firms Benchmark No p firms
Unemployment rate
Black-White gap -0.6731 0.0000 0.3009 0.0000
Average wage
Black 1.0749 1.0131 1.0755 1.0139
White 0.9984 1.0115 0.9985 1.0116
Wealth ratio
Wealth ratio 0.2028 0.2054 0.1611 0.1466
Consumption ratio
Consumption ratio 0.9717 0.9695 0.1861 0.1659
Note: The top panel compares the black-white unemployment rate, wage, consumption, and wealth gaps over the business cycles between the benchmark model and model without $p$ firms.
We further examine the behaviors in the cyclical properties of the average black-to-white wealth ratio and the average black-to-white consumption ratio. Other than the cyclicality of the average wealth ratio, the cyclical properties of the racial gaps in all other dimensions are reduced after discriminatory hiring is removed.
We further examine the cyclical properties of the average black-to-white wealth ratio and consumption ratio. In the benchmark model, the cyclicality of the average wealth ratio is 0.2028, which changes slightly to 0.2054 without $p$ firms, while its volatility decreases from 0.1611 to 0.1466. For the average consumption ratio, the cyclicality decreases marginally from 0.9717 to 0.9695, and the volatility drops more significantly from 0.1861 to 0.1659. These results show that, aside from the cyclicality of the wealth ratio, the cyclical properties of racial gaps across all other dimensions are notably reduced when discriminatory hiring is eliminated.
In summary, $p$ firms drive racial disparities in the labor market over the business cycle. The black-white unemployment rate gap, countercyclical in the benchmark, vanishes without $p$ firms. Wage cyclicality and volatility for black workers also decline, converging toward those of white workers. While the wealth ratio’s cyclicality rises slightly, its volatility and that of the consumption ratio decrease. These findings indicate that hiring discrimination is a key factor shaping racial differences in the cyclical properties of the labor market and economic outcomes.
5.4 Welfare disparities in business cycles
Lastly, we examine the business cycle welfare implications of removing hiring discrimination. We simulate the economy for 10,000 periods and compute the long-run average welfare for each individual type, following Cho and Ma (2023). The welfare measure is calculated as average consumption equivalence described in Section 4.3 over the simulation periods. We present the results in Table 8.
Table 8: Heterogeneous welfare change in business cycles
Average welfare gain (%) Eliminate $p$ firms
Overall -0.01
Black White
Average 0.02 -0.01
by productivity
Low 0.00 -0.01
Mid 0.02 -0.01
High 0.02 -0.02
by wealth
Low 20% 0.01 -0.01
40-60% 0.02 -0.02
Top 20% 0.03 -0.02
Note: This table compares the average consumption equivalence change from the benchmark model to the one without $p$ firms in business cycles. We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type. Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The economy experiences an average of 0.01% welfare decrease after removing prejudiced firms. This is much less than the welfare change in the steady state in Table 4, primarily because workers are risk averse. All workers experience a decrease in unemployment rate volatility, and white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions. On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%. Among white workers, those with high productivity and those in the top 20% of the wealth distribution lose the most (0.02%). In contrast, black workers in the top wealth quintile benefit the most, with an average welfare gain of 0.03%.
6 Conclusion
This study examines the effects of racial hiring discrimination within a frictional labor market on employment, wage disparities, and wealth accumulation between black and white workers in the U.S. We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation. Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in steady state and business cycles. Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum. Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome within frictional markets. Eliminating these discriminatory factors results in a decrease in overall welfare. Our findings shed light on the enduring nature of Black-White racial disparities in the U.S., offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
Appendix A Stationary Equilibrium
A stationary equilibrium consists of
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
    Unemployment insurance tax rate $\tau$
  6. 6.
    Dividends $d$
  7. 7.
    Distribution over employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and wealth ($a$), given by $\mu(e,R,s,a)$
such that:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
    Aggregate demand for capital equals aggregate supply, which in turn determines the interest rate $r$.
  5. 5.
[…]
  6. 6.
    Unemployment insurance tax rate $\tau$ solves to balance the government budget.
  7. 7.
    Dividend $d$ is the total flow profits of producing firms, net of total posting costs of vacant firms.
[…]
    where $j_{np}$ and $j_{p}$ refer to the flow profits of $np$ and $p$ firms respectively.
    $j_{np}(R,s,a)$ $=sf(k_{np})-(r+\delta)k_{np}-\omega_{np}$ (A.2)
    $j_{p}(wh,s,a)$ $=sf(k_{p})-(r+\delta)k_{p}-\omega_{p}$
  8. 8.
    The distribution $\mu(e,R,s,a)$ is invariant and is consistent with the optimal decision rules of capital choice, the law of motion of idiosyncratic productivity, and the labor market flows.
Appendix B Comparing the impact of bargaining power and extreme wealth shock
Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks. Equalizing black workers’ bargaining power to white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes. This is because non-prejudiced firms retain less profit, hence posting fewer vacancies. On the net, it also reduces white worker’s job-finding rate and average wage rate. As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
Table B.1: Aggregate impact of racial disparities from non-market factors
Moments Benchmark $\xi_{bl}=\xi_{wh}$ $\epsilon_{bl}=\epsilon_{wh}$
Households
job separation rate - black 0.06 0.06 0.06
job separation rate - white 0.04 0.04 0.04
job finding rate - black 0.50 0.43 0.48
job finding rate - white 0.66 0.64 0.66
unemp rate - black 0.12 0.13 0.12
unemp rate - white 0.05 0.05 0.05
mean wage - black 1.60 1.71 1.66
mean wage - white 2.13 2.12 2.15
mean wage ratio 0.75 0.81 0.77
mean wealth ratio 0.28 0.29 0.77
median wealth ratio 0.33 0.33 0.88
Firms
$p$ firm profit 0.04 0.04 0.04
np firm profit 0.03 0.02 0.02
p firm vacancy 0.02 0.02 0.02
np firm vacancy 0.05 0.04 0.05
p firm employment 0.54 0.58 0.56
np firm employment 0.40 0.35 0.38
Labor Market
$p$ market tightness 0.40 0.46 0.43
$np$ market tightness 0.73 0.60 0.69
Aggregate Outcomes
Y 3.06 3.05 3.08
K/Y 10.26 10.25 10.45
average wage 2.04 2.05 2.06
unemp rate 0.07 0.07 0.07
Note: This table compares the benchmark steady state to a model with equal bargaining power $\xi_{bl}=\xi_{wh}$ and a model with equal extreme wealth shock $\epsilon_{bl}=\epsilon_{wh}$.
In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers. The effect resembles assigning a higher bargaining power to black workers. This is because of the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012). With a lower probability of losing their wealth, black workers can accumulate more personal wealth. Higher personal wealth gives black workers higher reservation value when bargaining with firms. Effectively, black workers can bargain for higher wage outcomes (1.66 compared to 1.60 in the benchmark model). Similar to directly assigning a higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent. The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output. It spills over to an increase in the average wage rate for all. Though it has a much less impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap. Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).
Appendix C Impulse responses
We follow Boppart et al. (2018) to simulate the business cycle moments of the model from the derivatives of the impulse response functions. For the impulse response functions to represent the numerical derivative, the magnitude of the MIT shock should be small. This method also requires the business cycle dynamics to be well approximated as a linear system. Figure C.1 presents the impulse response functions for the benchmark model to a 1% positive and negative TFP shock at date 0. The shock gradually returns to a steady state with a persistence of 0.95. Our model produces symmetric impulse responses across all variables.
Figure C.1: Comparing impulse responses
Note: This figure plots the impulse response functions of the benchmark economy to a one percent increase and a one percent decrease in aggregate TFP at date zero. The red solid line denotes the responses to the expansionary shock, and the blue dashed line denotes the responses to the recessionary shock.

### v2

@@
 Abstract
-This paper examines the impact of racial discrimination in hiring on employment, wage, and wealth disparities between black and white workers.
+This paper examines the impact of racial discrimination in hiring on employment, wages, and wealth disparities between black and white workers.
 Using a labor search-and-matching model with racially prejudiced and non-prejudiced firms, we show that labor market frictions sustain discriminatory practices as an equilibrium outcome.
@@
 Discriminatory hiring also amplifies unemployment and wage volatility for black workers, increasing their labor market risks over business cycles.
-Eliminating prejudiced firms reduces these disparities and improves black workers’ welfare, though it slightly decreases overall economic welfare.
+Eliminating prejudiced firms reduces these disparities and improves the welfare of black workers, though it slightly decreases overall economic welfare.
 JEL classification: D14, E21, J15, J64, J65
@@
 Market competition and better information are expected to eliminate discrimination over time.
-However, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities.
+Nonetheless, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities.
 [^2] For example, Couch and Fairlie (2010); Biddle and Hamermesh (2013); Kuhn et al.
@@
 Small and Pager (2020) argues that discrimination reinforces itself across domains and constitutes a form of market failure.
-Only recently has macroeconomic research begun to explore racial disparities.
+Only recently has macroeconomic research begun exploring racial disparities.
 [^3] See, for example, Nakajima (2021), Aliprantis et al.
@@
 (2021).
-Yet, the role of discriminatory practices in shaping these disparities remains underexplored.
+Yet, the role of discriminatory practices in shaping these disparities remains under-explored.
 Developing a general equilibrium framework to explain the persistence of discrimination and its consequences is crucial for understanding the macroeconomic impacts of racial disparities and informing effective policymaking.
-This paper develops a labor search-and-matching framework to capture the persistence of hiring discrimination based on race.
+This paper constructs a labor search-and-matching framework to capture the persistence of hiring discrimination based on race.
 We show that hiring discrimination accounts for 44% to 52% of the black-to-white average wage gap and 16% of the racial median wealth gap.
-Additionally, hiring discrimination intensifies unemployment and wage volatility for black workers, creating a countercyclical racial unemployment gap and diverging racial wage cyclical patterns.
+Additionally, hiring discrimination increases both unemployment and wage volatility among black workers.
+It creates a countercyclical racial unemployment gap and diverges cyclical wage patterns between races.
 The model incorporates two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally.
-By construction, racial differences in the model are reflected as differences in bargaining power and the likelihood of experiencing wealth destruction.
+We include additional racial differences in the model as differences in bargaining power and the likelihood of experiencing wealth destruction.
 While all workers compete for jobs, white workers have additional opportunities with prejudiced firms.
@@
 Upon calibration, the model endogenously produces lower job-finding rates, higher unemployment rates, and lower bargained wages for black workers.
-Even without additional financial frictions, the model generates significant racial disparities in wealth accumulation.
+Even without further financial frictions, the model generates significant racial disparities in wealth accumulation.
 The main message of this paper is that labor market friction can sustain hiring discrimination as an equilibrium outcome.
@@
 In a counterfactual economy without prejudiced firms, the racial unemployment gap disappears, and the wage gap decreases significantly.
+Such a reduction in labor market disparity leads to asymmetric welfare changes.
 Black workers’ employment opportunities and wages increase because non-prejudiced firms expand their vacancy postings to fill positions left by prejudiced firms.
-As a result, black workers experience a welfare increase, with the largest gain among the high-productivity black workers.
+As a result, black workers experience a welfare increase, with the most significant gains among high-productivity black workers.
 White workers, however, lose the exclusive job opportunities that prejudiced firms previously provided.
-Their average wage declines and their unemployment rate rises.
+Their average wage declines, and their unemployment rate rises.
 This leads to a moderate welfare decrease, especially among the lowest-productivity and lowest-wealth white workers.
@@
 However, if we maintain the race-specific job destruction shocks, eliminating prejudiced hiring has a much smaller adverse impact on white workers.
-We further introduce aggregate fluctuations to the economy and show that hiring discrimination leads to distinct business cycle dynamics for black and white workers.
+We further introduce aggregate shocks to the economy and show that hiring discrimination generates distinct business cycle dynamics for black and white workers.
 Our model successfully replicates the observed countercyclical and volatile racial unemployment rate gap, as well as the more procyclical and volatile average wage rate for black workers.
@@
 (2017) that firm-side discrimination contributes to the business cycle differences between black and white workers, and non-market demographic factors alone do not fully account.
-Lastly, we examine the business cycle welfare change between black and white workers when discriminatory hiring contributes to more volatile labor market outcomes for black workers.
-We compare the average consumption equivalence of workers during business cycles after removing prejudiced firms.
-Not all workers benefit from a less volatile economy without discriminatory firms.
+Lastly, we examine the associated business cycle welfare change between black and white workers when discriminatory hiring contributes to more volatile labor market outcomes for black workers.
+Not all workers benefit from a less volatile economy without prejudiced firms.
 Black workers gain an average of 0.02% in welfare, whereas white workers face a welfare loss of 0.01%.
@@
 While much of the literature focuses on static disparities, fewer studies document how macroeconomic fluctuations impact black and white workers differently.
-Notably, Couch and Fairlie (2010) demonstrate that black workers are disproportionately affected during economic downturns, being the last hired in upturns and the first fired in downturns.
+Notably, Couch and Fairlie (2010) demonstrate that black workers are disproportionately affected during recessions, being the last hired in upturns and the first fired in downturns.
 Biddle and Hamermesh (2013) find that the discriminatory wage gap is procyclical, widening during economic expansions.
@@
 For example, Derenoncourt et al.
-(2023) provide a historical account of wealth segregation over the past 150 years, showing persistent and substantial disparities.
+(2023) provides a historical account of wealth segregation over the past 150 years, showing persistent and substantial disparities.
 Boerma and Karabarbounis (2021) and Aliprantis et al.
@@
 Furthermore, Ganong et al.
-(2020) demonstrate that income risks affect individuals differently across racial groups due to wealth differences.
+(2020) demonstrates that income risks affect individuals differently across racial groups due to wealth differences.
 Building on these findings, Bartscher et al.
@@
 Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over into long-term wealth accumulation disparities between black and white workers.
-Germane to our project, Nakajima (2021) develop a search-and-matching model to examine how monetary policies exacerbate racial differences in the labor market.
-Our model explicitly distinguishes between discriminatory and non-discriminatory firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy.
+Germane to our project, Nakajima (2021) develops a search-and-matching model to examine how monetary policies exacerbate racial differences in the labor market.
+Our model explicitly distinguishes between prejudiced and non-prejudiced firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy.
 We analyze the spillover effects of discriminatory hiring practices on both labor and wealth disparities, offering a comprehensive theoretical understanding of these dynamics over the long run and across business cycles.
@@
 Borella et al.
-(2018) show that incorporating gender differences in a life-cycle model improves its empirical fit, emphasizing the importance of accounting for demographic heterogeneity.
-Krusell and Smith (1998) conclude that heterogeneity in wealth does not alter business cycle fluctuations.
+(2018) shows that incorporating gender differences in a life-cycle model improves its empirical fit, emphasizing the importance of accounting for demographic heterogeneity.
+Krusell and Smith (1998) concludes that heterogeneity in wealth does not alter business cycle fluctuations.
 Yet, Yum et al.
-(2023) show that heterogeneity could generate large aggregate fluctuations when introducing non-convexity in budget constraint through progressive tax.
+(2023) shows that heterogeneity could generate large aggregate fluctuations when introducing non-convexity in budget constraints through progressive tax.
 Our paper demonstrates that incorporating racial differences into labor market models enhances their ability to capture business cycle fluctuations.
@@
 The rest of the paper proceeds as follows.
-Section 2 lays out theoretical framework.
+Section 2 lays out the theoretical framework.
 Section 3 discusses the calibration strategy.
 Section 4 examines the steady-state implications of racial discrimination.
-Section 5 provides business cycle implications.
+Section 5 provides the business cycle implications.
 Section 6 concludes the paper.
 2 Model
-In this section, we present a model of labor market discrimination without aggregate uncertainty.
+This section presents a model of labor market discrimination without aggregate uncertainty.
 We use this framework to understand the role of hiring discrimination in driving the racial gaps in labor market outcomes and wealth.
@@
 Non-prejudiced firms ($np$) hire both black and white workers without discrimination.
-Prejudiced and non-prejudiced firms also have type-specific job destruction shocks.
+Prejudiced and non-prejudiced firms also experience type-specific job destruction shocks.
+Each firm posts one vacancy or hires one worker.
+Thus, firms and jobs are synonymous in our model.
 To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote the variables in the next period.
@@
 The workers face idiosyncratic productivity shocks $s$, following an AR(1) process $\log(s^{\prime})=\rho_{s}\log(s)+\epsilon_{s}$, with $\epsilon_{s}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{s}^{2})$.
-Workers also differ in terms of their asset holdings.
-They have access to risk-free assets, and the workers can save using these assets to partially insure themselves against the labor market risks.
+Workers also differ in their asset holdings.
+They can save using these risk-free assets to partially insure themselves against the labor market risks.
 Following Mukoyama (2013), workers also receive race-specific extreme wealth shocks with probability $\epsilon_{R}\in\{\epsilon_{bl},\epsilon_{wh}\}$.
 Upon realization of the shock, the worker loses all their wealth.
-Taken together, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$), and the endogenous distribution of workers is given by $\mu(e,R,s,a)$.
+Altogether, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$).
+The endogenous distribution of workers is summarized by $\mu(e,R,s,a)$.
 All endogenous aggregate variables depend on $\mu$.
 Individual workers may move around the distribution between periods; however, the overall distribution remains stationary.
-We omit the record of $\mu$ in the steady state and will reintroduce $\mu$ when augmenting the model with aggregate shock.
+We omit the record of $\mu$ in the steady state and will reintroduce $\mu$ when augmenting the model with aggregate shocks in Section 5.
 2.2 Labor market search and matching
-The total number of unemployed workers $u$ is the sum of unemployed black ($u_{bl}$) and white ($u_{wh}$) workers.
-The number of $np$ vacancies available is $v_{np}$, while the number of $p$ firm vacancies is $v_{p}$.
-Since the non-prejudiced firms search among both black and white unemployed workers, the non-prejudiced market tightness is given by $\theta_{np}=v_{np}/u$.
-On the other hand, prejudiced firms employ only white unemployed workers, and hence, the prejudiced market tightness is $\theta_{p}=v_{p}/u_{wh}$.
-Following Den Haan et al.
+The total number of unemployed workers, denoted as $u$, is the sum of unemployed black workers $(u_{bl})$ and unemployed white workers $(u_{wh})$.
+There are $v_{np}$ vacancies available in the non-prejudiced market, while the number of vacancies in the prejudiced market is $v_{p}$.
+Non-prejudiced firms search among both black and white unemployed workers, resulting in a non-prejudiced market tightness defined as $\theta_{np}=\frac{v_{np}}{u}$.
+In contrast, prejudiced firms only hire unemployed white workers, which gives the prejudiced market tightness $\theta_{p}=\frac{v_{p}}{u_{wh}}$.
+Following the works of Den Haan et al.
 (2000) and Petrosky-Nadeau et al.
-(2018), unemployed workers and vacant firms match via a constant return to scale matching function
+(2018), unemployed workers and vacant firms match through a constant returns to scale matching function
 $M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}},$ (1)
@@
 2.3 Unemployment Insurance
-Unemployment insurance is characterized by the replacement rate $h$, probability of eligibility $P_{e}$, and maximum eligibility level $\chi$.
+Unemployment insurance is characterized by the replacement rate $h$, eligibility probability $P_{e}$, and maximum coverage level $\chi$.
 Following Setty and Yedid-Levi (2021), eligible workers receive unemployment benefits $b(R,s,a)=min\{h\bar{w}(R,s,a),\chi\}$, where $\bar{w}(R,s,a)$ is the counterfactual wage earned by an employed worker with race $R$, productivity $s$, and wealth $a$.
@@
 Similar to Mitman and Rabinovich (2015), newly unemployed workers receive unemployment benefits with certainty and continue to receive benefits next period with probability $P_{e}$.
-If an unemployed worker loses their eligibility to receive benefits, they continue to remain ineligible in the future as well.
-Unemployment benefits are funded through proportional tax $\tau$ on the labor income, and the government sets $\tau$ to balance its budget.
+If an unemployed worker loses their eligibility to receive benefits, they continue to remain ineligible in the future.
+Unemployment benefits are funded through a proportional tax $\tau$ on the labor income, and the government sets $\tau$ to balance its budget.
 [^4] We intentionally model a more realistic and complex unemployment insurance structure to capture the racial disparities in the incidence and take-up of unemployment insurance.
@@
 Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced ($U^{I}_{np}(R,s,a)$) or a prejudiced ($U^{I}_{p}(wh,s,a)$) firm previously.
-An unemployed worker who is not eligible for unemployment benefits earns $U^{N}(R,s,a)$ over their lifetime.
+An unemployed worker who is not eligible for unemployment benefits has $U^{N}(R,s,a)$ over their lifetime.
 All the workers discount their future utility by $\beta$.
@@
 2.4.1 Employed with np firm
-Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in an $np$ firm.
-The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their lifetime discounted utility.
+Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in a $np$ firm.
+The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their discounted lifetime utility.
 Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$.
@@
 The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3.
-$\hat{W}_{np}(R,s^{\prime},a^{\prime})=\lambda_{np}U^{I}_{np}(R,s^{\prime},a^{\prime})+(1-\lambda_{np})W_{np}(R,s^{\prime},a^{\prime})$ (3)
 If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$.
@@
 If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
+$\hat{W}_{np}(R,s^{\prime},a^{\prime})=\lambda_{np}U^{I}_{np}(R,s^{\prime},a^{\prime})+(1-\lambda_{np})W_{np}(R,s^{\prime},a^{\prime})$ (3)
 2.4.2 Employed with p firm
-Prejudiced firms $p$ hire only white workers. The rest of the problem is similar to the previous case and is given by Equation 4.
+Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, and we only keep track of white workers since prejudiced firms do not hire black workers.
 $W_{p}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{W}_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{W}_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (4)
@@
 2.4.3 Unemployed and eligible workers
-Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or an $np$ firm.
+Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or a $np$ firm.
 In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
-Black worker from a $np$ firm
+Unemployed black worker from $np$ firm
 Equation 6 describes the value of an unemployed black worker eligible for benefits.
-The worker receives unemployment insurance payout, $b_{np}(bl,s,a)$, net of labor income tax rate $\tau$, previous savings and lump-sum dividend transfer $d$.
 The worker encounters the extreme wealth shock with probability $\epsilon_{bl}$.
+The worker receives unemployment insurance payout, $b_{np}(bl,s,a)$, net of labor income tax rate $\tau$, previous savings, and lump-sum dividend transfer $d$.
 $U^{I}_{np}(bl,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{bl})\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})+\epsilon_{bl}\hat{U^{I}}_{np}(bl,s^{\prime},0)\big{]}\Big{\}}$ (6)
@@
 $\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))[P_{e}U^{I}_{np}(bl,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(bl,s^{\prime},a^{\prime})]$ (7)
-The unemployed worker finds a job with probability $f(\theta_{np})$.
-If the worker doesn’t get matched, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $1-P_{e}$.
-White worker from a $np$ firm
+The unemployed worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$.
+Otherwise, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$.
+If one loses the benefit, their value function becomes $U^{N}(bl,s^{\prime},a^{\prime})$.
+Unemployed white worker from $np$ firm
 Unemployed white workers are recruited by both $p$ and $np$ firms.
-Their value function is described by Equation 8.
-Like a black worker, a white worker has cash-on-hand from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer.
-They also encounter an extreme wealth shock with probability $\epsilon_{wh}$.
+Equation 8 describes their value function.
+Like a black worker, a white worker encounters an extreme wealth shock with probability $\epsilon_{wh}$.
+They also have a budget constraint from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer.
 $U^{I}_{np}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U^{I}}_{np}(wh,s^{\prime},0)\big{]}\Big{\}}$ (8)
@@
 Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$.
-They can receive a job offer from a $p$ firm with probability $(1-f(\theta_{np}))f(\theta_{p})$, an offer from an $np$ firm with probability $f(\theta_{np})(1-f(\theta_{p}))$, and receive both offers together with probability $f(\theta_{np})f(\theta_{p})$.
-The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$) and continue to receive unemployment benefits with probability $P_{e}$.
-If the worker receives both $p$ and $np$ offers, they will choose the job with higher lifetime utility.
-$\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (9)
-$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
-$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{np}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
-White worker from a $p$ firm
+With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives an only job offer from a $p$ firm.
+With probability $f(\theta_{np})(1-f(\theta_{p}))$, they receive an only offer from a $np$ firm.
+With probability $f(\theta_{np})f(\theta_{p})$, they receive offers from both $np$ and $p$ firms simultaneously.
+If the worker receives both $p$ and $np$ offers, they will choose the job with a higher lifetime value.
+The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$).
+In this case, one may continue to receive unemployment benefits with probability $P_{e}$.
+Otherwise, one moves to a state of receiving no unemployment benefit with probability $(1-P_{e})$.
+$\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (9)
+$+f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
+$+(1-f(\theta_{p}))(1-f(\theta_{np}))\Big{[}P_{e}U^{I}_{np}(wh,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})\Big{]}$
+Unemployed white worker from $p$ firm
 Equation 10 describes the problem faced by a white worker who last worked with a $p$ firm.
@@
 $c+a^{\prime}$ $=(1-\tau)b_{p}(wh,s,a)+(1+r)a+d$
-After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the worker’s future value is described in Equation 11.
-It is nearly identical to Equation 9 except for the value of remaining unemployed from a $p$ firm rather than a $np$ firm.
-$\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (11)
-$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
-$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
+After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the unemployed worker’s future value is described in Equation 11. It is analogous to Equation 9.
+$\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (11)
+$+f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
+$+(1-f(\theta_{p}))(1-f(\theta_{np}))\Big{[}P_{e}U^{I}_{p}(wh,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})\Big{]}$
 2.4.4 Unemployed and ineligible workers
@@
 They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again.
-Thus, their income comes only from their previous savings and lump-sum dividends and does not depend on their prior employment situation.
+Therefore, their consumption relies solely on previous savings and lump-sum dividends, independent of their prior employment status.
 Black worker
-Equation 12 describes the problem of a black worker searching for jobs without unemployment insurance.
+Equation 12 describes the problem of an unemployed black worker searching for jobs without unemployment insurance.
 $U^{N}(bl,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{bl})\hat{U^{N}}(bl,s^{\prime},a^{\prime})+\epsilon_{bl}\hat{U^{N}}(bl,s^{\prime},0)\Big{]}\Big{\}}$ (12)
@@
 $c+a^{\prime}$ $=(1+r)a+d$
-And Equation 13 expands the future value of the black worker.
+And Equation 13 expands the future value of the black worker. One has a probability of $f(\theta_{np})$ of finding a job - only with a $np$ firm; otherwise, they remain unemployed and ineligible.
 $\hat{U^{N}}(bl,s^{\prime},a^{\prime})$ $=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))U^{N}(bl,s^{\prime},a^{\prime})$ (13)
@@
 And Equation 15 expands the future value of the white worker in this scenario.
-$\hat{U^{N}}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (15)
-$+\underbrace{f(\theta_{p})f(\theta_{np})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth np and p sector jobs}}$
-$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing unemployed}}$
-Thus, the problem faced by an ineligible worker is analogous to that faced by an eligible worker, except that they don’t receive unemployment benefits.
+It is nearly identical to Equation 11.
+With a probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$, the unemployed worker does not find any job and remains unemployed and ineligible.
+$\hat{U^{N}}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (15)
+$+f(\theta_{p})f(\theta_{np})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
+$+(1-f(\theta_{p}))(1-f(\theta_{np}))U^{N}(wh,s^{\prime},a^{\prime})$
 2.5 Firms
-There are a large number of $p$ and $np$ firms that post vacancies after paying the vacancy posting costs.
-The firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
+A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. These firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
 2.5.1 Vacant np firm
@@
 Equation 16 describes the value of a vacant $np$ firm, $V_{np}$.
-$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}}_{\text{value of matching with atype (s,a) black worker}}\Big{]}$ (16)
-$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{np provides higher value than a p firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}}_{\text{value of matching with (s,a) white worker}}\Big{)}$
-$+\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{p provides higher value than a np firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))}_{\text{value of matching (s,a) worker without competing p offer}}\Big{]}\Big{\}}da,$
+$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}\Big{]}$ (16)
+$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}$
+$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))\Big{]}\Big{\}}da,$
 The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$.
@@
 These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$.
-Since the white unemployed workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ match and begin producing only if it gives them a higher value or if the worker did not receive a $p$ firm offer.
-The last two rows of Equation 16 reflect the cases that a white worker will produce with the $np$ firm.
+Since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}})$, or if the worker did not receive a $p$ firm offer, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}(1-f(\theta_{p})))$.
 2.5.2 Vacant p firm
 Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers.
-The probability of matching with a white unemployed worker is $q(\theta_{p})$.
+The probability of matching with an unemployed white worker is $q(\theta_{p})$.
 Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
-$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}}_{\text{value of matching with (s,a) white worker, $p$ offers higher value than a np firm}}$ (17)
-$+\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))}_{\text{value of matching with (s,a) white worker without competing np offer}}\Big{]}da$
-Since the vacant $p$ firm searches only among white workers, the probability of matching with an unemployed white worker having productivity $s$ and asset $a$ is given by $\phi_{u}(wh,s,a)/u_{wh}$.
-Like the case of $np$ firms, the white worker accepts the $p$ match only if the $p$ offer is more favorable or if they don’t receive a competing $np$ offer.
+$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}$ (17)
+$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))\Big{]}da$
+Since the vacant $p$ firm only searches among white workers, the probability of matching with an unemployed white worker who has productivity $s$ and assets $a$ is represented by $\phi_{u}(wh,s,a)/u_{wh}$.
+Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than other options or if they do not receive a competing offer from an $np$ firm.
 We assume that there is free entry of firms, and hence both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
@@
 The value of a producing $np$ firm is described by Equation 18.
-$j(R,s,a)$ describes the period profit.
-The worker’s productivity $s$ enters the production function.
-The matched firm produces $sf(k)$ units of output, pays the rental and the depreciation cost of capital, $(r+\delta)k$, and the wage cost, $\omega_{np}(R,s,a)$.
-$J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\Big{(}\frac{1-\lambda^{np}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
-where
-$j(R,s,a)$ $=sf(k)-(r+\delta)k-\omega_{np}(R,s,a).$
+The firm chooses capital $k$ to maximize its lifetime value producing with the worker.
 The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$.
@@
 The future value of the firm incorporates this probability.
+In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output minus production costs.
+The worker’s productivity, $s$, is a factor in the production function $sf(k)$.
+The matched firm produces $sf(k)$ units of output and incurs costs for capital rental and depreciation, represented by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
+$J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\frac{1-\lambda^{np}}{1+r}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
+where
+$j(R,s,a)$ $=sf(k)-(r+\delta)k-\omega_{np}(R,s,a).$
 2.6 Producing p firm
 Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm.
-The firm hires capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
+The firm rents capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
 $J_{p}(wh,s,a)$ $=\max_{k}\Big{\{}j(wh,s,a)+\Big{(}\frac{1-\lambda^{p}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})J_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}J_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (19)
@@
 A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$.
-The worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$.
+A worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$.
 The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
@@
 $\omega_{p}(wh,s,a)$ $=\operatorname*{arg\,max}_{\omega_{p}}(W_{p}(wh,s,a)-U^{I}_{p}(wh,s,a))^{\xi_{wh}}J_{p}(wh,s,a)^{1-\xi_{wh}}$ (21)
-We define the stationary equilibrium in Appendix A.
+Finally, we define the stationary equilibrium in Appendix A.
 3 Calibration
-A period in the model represents a quarter.
+Each period in the model represents a quarter.
 We calibrate the benchmark model to match the relevant U.S.
@@
 We have two sets of parameters.
-The first group of parameters is chosen externally based on literature and empirical evidence without using model-generated data, while the second group of parameters is calibrated internally by simulating our model to match a set of relevant data moments.
+One group of parameters is chosen externally based on literature and empirical evidence without using model-generated data.
+The other parameters are calibrated internally by simulating our model to match a set of relevant data moments.
 We set the share of black workers to be 19.1% in the model, with the remaining being white workers.
-Table 1 shows the internally calibrated parameter values and their targeted moments, as well as the externally chosen parameters and their sources.
+Table 1 shows the internally calibrated parameter values, their targeted moments, the externally chosen parameters, and their sources.
 Table 1: Calibration and targeted statistics
@@
 $\kappa_{p}$ 4.2622 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
-$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1 1
+$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1.0 1.0
 $\lambda_{np}$ 0.0644 np sector job destruction shock job separation rate - black 0.0644 0.0644
@@
 Chosen externally
-$\alpha$ 0.6600 elasticity of labor matching Nakajima (2012)
-$\theta_{n}$ 0.2890 capital share of output Nakajima (2012)
+$\alpha_{k}$ 0.2890 capital share of output Nakajima (2012)
 $\delta$ 0.0150 quarterly depreciation rate Nakajima (2012)
@@
 $h$ 0.4000 UI replacement rate Mitman and Rabinovich (2015)
-$\xi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
+$\chi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
 $Pe$ 0.5385 probability of UI eligibility maximum weeks of eligibility
 Notes: This table reports the parameters, their values, and descriptions.
-The top panel presents the parameters chosen internally by minimizing the distance between model generated moments and data.
+The top panel presents the parameters chosen internally by minimizing the distance between model-generated moments and data.
 The last two columns of the top panel compare the targeted moments in the data and in the model simulations.
@@
 3.2 Production
-The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha}$.
-We choose $\alpha$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015 following Nakajima (2012).
+The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$.
+We choose $\alpha_{k}$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015, following Nakajima (2012).
 3.3 Productivity and wealth shocks
@@
 Our estimation strategy closely follows Setty and Yedid-Levi (2021).
-We run a standard Mincer wage regression of log wage on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of males aged 25 years and above.
+We run a standard Mincer wage regression of log wages on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of men 25 years and older.
 We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data.
-The estimated regression is at a biennial frequency, given that the PSID data is available once in two years.
+The estimated regression is at a biennial frequency, given that the PSID data is available once every two years.
 We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process.
 The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency.
-We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.
+We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.555We follow Setty and Yedid-Levi (2021) to have only men in the sample.
+We also calculate the AR(1) processes, including men and women from PSID.
+The resulting $\rho_{s}$ is unchanged from the men-only sample (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
 Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth.
 The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers.
-They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth.
-(Nakajima, 2021).
+They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth (Nakajima, 2021).
 3.4 Unemployment insurance
 The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$.
-We choose the replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others.
+We choose the unemployment insurance replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others.
 Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model.
@@
 (2017).
-The elasticity of the matching function, $\iota$, targets the job finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al.
+The elasticity of the matching function, $\iota$, targets the job-finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al.
 (2000) and Petrosky-Nadeau et al.
 (2018).
-We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers respectively.
+We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively.
 The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively.
 [^6] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers.
-On the other hand, white workers work in both $p$ and $np$ firms, and hence $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380).
-The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, to be 1 following Wolcott (2021).
+On the other hand, white workers work in both $p$ and $np$ firms; hence, $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380).
+The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, of 1.0, following Wolcott (2021).
 [^7] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$.
@@
 The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012).
-Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.211.
+Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.2110.
 3.6 Assessing the model as a quantitative theory of racial disparity
-Our calibration successfully captures the racial gaps along three important dimensions: income, wealth, and labor market outcomes.
-We now discuss each of these dimensions in more detail before proceeding with the quantitative exercises examining the macroeconomic impact of racial discrimination.
+Our calibration successfully captures the racial gaps across three important dimensions: income, wealth, and employment outcomes.
+We now discuss these dimensions in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
 (a) Labor income share
@@
 Figure 1: Labor Income and Wealth Distribution
-Note: This figure compares the steady state model generated distributions of wealth and labor income with their empirical analogs.
-The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Survey of Consumer Finance (SCF), respectively.
-First, the model successfully captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we don’t target these distributions explicitly.
-As with the majority of incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile.
+Note: This figure compares the model-generated steady-state distributions of wealth and labor income with their empirical analogs.
+The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF).
+First, the model captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not target these distributions explicitly.
+As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile.
 On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
 Note: This table compares the steady state wealth and unemployment moments between black and white workers with the corresponding empirical data.
-Second, the model is able to capture the disparities in unemployment, labor income, and wealth between black and white workers.
+Second, the model reproduces the disparities in unemployment, labor income, and wealth between black and white workers.
 As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth.
-In addition to the targeted moments, Table 2 demonstrates that our model successfully captures the untargeted employment and wealth moments.
-Specifically, our model reproduces the 5% unemployment rate among white workers compared to 12% among black workers.
-Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to the prejudiced sector.
-Our model results show that the $p$ sector with the equilibrium market tightness of 0.4, though smaller than the $np$ market tightness of 0.73, provides an important advantage for white workers by increasing their job-finding rate compared to black workers.
+In addition to the targeted moments, Table 2 demonstrates that our model successfully generates the untargeted employment and wealth moments.
+Specifically, our model produces a 5% unemployment rate among white workers compared to 12% among black workers.
+Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to prejudiced firms.
 On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al.
@@
 However, the model understates the median wealth gap between black and white workers.
-The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in the data (Kaplan et al., 2014).
-In sum, the model captures the racial inequalities across labor income, wealth, and unemployment outcomes. We next explore how racial discrimination in hiring contributes to these differences.
+The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in Kaplan et al.
+(2014).
+In summary, the model reflects the racial disparities present in labor income, wealth, and unemployment outcomes.
+Next, we will investigate how racial discrimination in hiring contributes to these disparities.
 4 Steady state results
@@
 4.1 Sustaining racial discrimination
-The labor search-and-matching framework allows firms to earn profit from the surplus generated by the search frictions.
+The labor search-and-matching framework allows firms to earn a profit from the surplus generated by search frictions.
 This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium.
-It contrasts with a frictionless canonical neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957).
-In particular, our benchmark model has $p$ firms retaining a profit of 0.04 while $np$ firms retain a profit of 0.03 in the steady state (see Column 1 of Table 3).
-Next, we examine the steady-state interplay between discriminatory and non-discriminatory firms.
-Since discriminatory hiring remains profitable in the benchmark equilibrium, we gradually increase penalties by raising the vacancy posting costs for discriminatory firms until they are no longer profitable.
-Figure 2 displays the dynamics of employment, vacancy, and profits of the prejudiced and non-prejudiced firms.
+It contrasts with a canonical frictionless neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957).
+Table 3 shows firm performances in the steady state.
+$p$ firms have a total profit of 0.04, while $np$ firms have 0.03.
+This difference arises because $p$ firms experience a lower job destruction shock (0.0268) than $np$ firms (0.0644).
+Although $p$ firms incur higher vacancy posting costs (4.2622) to hire only white workers, they retain workers longer (employment level at 0.54) at lower posting rates (0.02).
+Conversely, $np$ firms face lower vacancy posting costs (2.735), but their high job destruction shock leads to a lower employment level (0.40) and a more frequent need for vacancy postings (0.05).
+Table 3: Firms in the steady state
+Moments Benchmark
+$p$ firm profit 0.04
+np firm profit 0.03
+p firm vacancy 0.02
+np firm vacancy 0.05
+p firm employment 0.54
+np firm employment 0.40
+Note: This table shows steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
+Next, we examine the steady-state interplay between prejudiced and non-prejudiced firms.
+At the benchmark equilibrium, $p$ firms post a positive number of vacancies with sufficient expected profits to cover the posting costs.
+We gradually increase penalties for them by raising $\kappa_{p}$ until their expected operating profit no longer justifies vacancy postings.
+Figure 2 displays the profit, vacancy, and employment dynamics between the $p$ and $np$ firms.
 As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment.
@@
 It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks.
-As $p$ firms’ profit declines, the profit of $np$ firms increases at a slower rate.
-Additionally, as $np$ firms replace $p$ firms, they recruit from both black and white workers, rather than only white workers displaced from $p$ firms.
+As $p$ firms’ profits decline, the profits of $np$ firms increase at a slower rate.
+Additionally, as $np$ firms replace $p$ firms, $np$ firms recruit from both black and white workers, rather than only white workers displaced from $p$ firms.
 This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
-In the limit, as $\kappa_{p}$ approaches the total expected value of the $p$ sector, discriminatory ($p$) firms almost completely exit the market, since posting vacancies exclusively for white workers becomes unprofitable.
-The value of $\kappa_{p}$ at which $p$ firms exit is 217% of the benchmark value of 4.26.
-Meanwhile, non-discriminatory ($np$) firms become nearly the sole employers in the economy.
-The gray dotted lines in Figure 2 illustrate a counterfactual steady state where we entirely remove $p$ firms from the benchmark model.
-Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of completely eliminating $p$ firms.
+In the limit, as $\kappa_{p}$ approaches the total expected value of an operating $p$ firm, posting vacancies exclusively for white workers becomes entirely unprofitable.
+As a result, all $p$ firms exit the market.
+The value of $\kappa_{p}$ at which $p$ firms exit entirely is 217% of the benchmark value of 4.2622.
+Meanwhile, $np$ firms become the sole employers in the economy.
+The gray dotted lines in Figure 2 illustrate a steady state where we construct a counterfactual model that structurally removes $p$ firms from the benchmark model.
+Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of structurally eliminating $p$ firms.
 Figure 2: Penalizing discrimination
@@
 4.2 Impact of hiring discrimination on racial inequality
-In this subsection, we compare the benchmark economy to one without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring.
+In this subsection, we compare the benchmark economy to those without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring.
 Table 4 presents the comparisons.
 All firms in the counterfactual models provide equal hiring opportunities to black and white workers.
-Column 2 removes the $p$ firms from the model, which leaves all workers with equal job destruction shocks at the benchmark $\lambda_{np}$.
-In Column 3, we re-calibrate the job destruction shocks to race-specific and have $\lambda_{wh}$ and $\lambda_{bl}$ as race-specific job destruction shocks.
+Column 2 removes the $p$ firms from the benchmark model, which leaves all workers with equal job-finding rates with the $np$ firms and equal job destruction shocks at the benchmark $\lambda_{np}$.
+In Column 3, we further adjust the job destruction shocks to be race-specific, with $\lambda_{wh}$ and $\lambda_{bl}$ designating the shocks for white and black workers, respectively.
 [^8] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates.
 This counterfactual economy only has one type of firm.
-We assign white workers a job destruction shock, $\lambda_{wh}$, to be 0.038 and black workers a destruction shock, $\lambda_{bl}$, to be 0.0644 to keep black and white workers’ job separation rates as in the calibration.
-Re-assigning the job destruction shocks further isolates the impact of job separation.
+We assign white workers a job destruction shock, $\lambda_{wh}$, of 0.0380 and black workers a destruction shock, $\lambda_{bl}$, of 0.0644 to maintain black and white workers’ job separation rates as in the calibration.
+Reassigning the job destruction shocks further isolates the impact of job separation.
 Table 4: Impact of hiring discrimination
@@
 All other parameters stay the same as benchmark calibration.
-The first panel shows the changes in racial employment, wage, and wealth disparities after removing prejudiced hiring.
+The first panel illustrates the changes in racial employment, wage, and wealth disparities after removing prejudiced firms.
 The first two rows describe the job separation rates.
-By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the difference as the benchmark.
+By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the differences as the benchmark calibration.
 The next two rows describe the job-finding rates.
-Since there is no $p$ sector in the economy in Columns 2 and 3, the job-finding rates are equalized for black and white workers.
-White workers only have one sector for employment.
-Hence, their job-finding rate reduces.
-Black workers have more job opportunities as the $np$ sector posts more vacancies; hence, their job-finding rate increases.
+Since there are no $p$ firms in the economy in Columns 2 and 3, the job-finding rates are equal for black and white workers.
+White workers have only one sector for employment; hence, their job-finding rate decreases.
+Black workers have more job opportunities as the $np$ sector posts more vacancies; thus, their job-finding rate increases.
 Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2.
-Altogether, it reduces the racial gap in the unemployment rate.
+Altogether, removing $p$ firms reduces the racial gap in the unemployment rate.
 In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes.
-In Column 3, the unemployment rate gap closes from 7% to 3% because of the remaining racial difference in $\lambda$.
-Turning to the wage impact, black workers can bargain for a higher wage rate as they receive more job opportunities when we remove discriminatory hiring.
-White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rate.
-As a result, the average black-to-white wage ratio increases from 75% in benchmark to 88% in Column 2 and to 86% in Column 3.
+In Column 3, the unemployment rate gap closes from 7% to 3% due to the remaining racial difference in $\lambda$.
+Turning to the wage impact, black workers can bargain for higher wage rates as they receive more job opportunities when we remove discriminatory hiring.
+White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rates.
+As a result, the average black-to-white wage ratio increases from 75% in the benchmark to 88% in Column 2 and to 86% in Column 3.
 We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
 Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44.
-The result remains stable between Columns 2 and 3.
+The results remain stable between Columns 2 and 3.
 We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
@@
 The second panel displays the changes in firm attributes.
-We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock.
-Without discriminatory firms, $np$ firms’ profit doubles, vacancy posting, and employment level increase.
+[^10] We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock.
+Without prejudiced firms, $np$ firms’ profits double, and vacancy postings and employment levels increase.
 These changes result in the aforementioned advantages for black workers in the counterfactual economies.
-In Column 3, white workers’ lower job destruction shock drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
+In Column 3, the lower job destruction shock for white workers further drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
 The third panel of Table 4 displays the labor market conditions.
-In response to the $p$ firms exiting, more $np$ firms endogenously enter the market.
-The labor market tightness of the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms and 1.12 when we split $\lambda$ by race.
+As $p$ firms exit, a greater number of $np$ firms enter the market endogenously.
+The labor market tightness for the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms, and to 1.12 when we further split $\lambda$ by race.
 The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
 The last panel of Table 4 describes the aggregate outcomes.
-With more $np$ firms entering, the aggregate output stays relatively unchanged in Column 3.
+With more $np$ firms entering, the aggregate output remains relatively unchanged in Column 3.
 The capital-output ratio slightly decreases to 10.25.
-The economy-wide average wage increases to 2.05 from benchmark 2.04, and the aggregate unemployment rate decreases to 6%.
-However, Column 2 experiences a more significant reduction in output, average wage, and higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs rather than a resource for production.
-Our results show that the presence of discriminatory firms can explain sizable fractions of income and wealth inequality.
-Penalizing prejudiced firms primarily improves the employment opportunities and wage outcomes for the black workers; it also worsens the situation for the white workers.
-Removing $p$ firms reduces the job-finding rate and, hence, the unemployment rate for white workers.
-In addition, they also lower the value of their outside option in wage bargaining, which, in turn, affects their income and wealth outcomes.
-To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by the black and the white workers.
+The economy-wide average wage increases to 2.05 from the benchmark of 2.04, and the aggregate unemployment rate decreases to 6%.
+However, Column 2 experiences a more significant reduction in output and average wage, along with a higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs instead of using resources for production.
+Our results show that the presence of prejudiced firms can explain sizable fractions of income and wealth inequality.
+Penalizing prejudiced firms primarily improves employment opportunities and wage outcomes for black workers; it also worsens the situation for white workers.
+Removing $p$ firms decreases job-finding rates and increases unemployment for white workers, lowering their bargaining positions and affecting their income and wealth.
+To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by black and white workers.
 4.3 Welfare analysis
@@
 We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change.
-We aggregate over the counterfactual distribution rather than benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms.
+We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms.
 Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift.
-We present the average welfare change between black and white workers by productivity types and by wealth quintiles in Table 4.
-Overall, the economy experiences a 5.8% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race.
+We present the average welfare changes for black and white workers by productivity types and wealth quintiles in Table 5.
+Overall, the economy experiences a 5.83% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race.
 Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%.
@@
 On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%.
-This is because of the decrease in their wage rate and the increase in the unemployment rate.
+This is because of the decrease in their wage rates and the increase in the unemployment rate.
 Mid-panels in Table 5 present the heterogeneous welfare change for black and white workers by productivity types.
-Though the welfare reduction of eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms.
+Though the welfare reduction from eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms.
 In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race.
-In comparison, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
+Furthermore, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
 Table 5: Heterogeneous welfare change
@@
 The bottom panels of Table 5 present the welfare change by wealth quintiles for black and white workers.
-Removing discriminatory firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race.
-The most considerable reduction in welfare (9.43%) happens for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
+Removing prejudiced firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race.
+The most considerable reduction in welfare (9.43%) occurs for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
 5 Business cycle dynamics
-In this section, we first document the racial disparities over the business cycle.
-Then, we introduce aggregate uncertainty in our benchmark model to quantify the role of hiring discrimination in generating disparate dynamics in the business cycle.
-Lastly, we discuss the welfare implications of hiring discrimination over the business cycle.
+In this section, we begin by documenting the racial disparities over business cycles.
+Next, we incorporate aggregate uncertainty into our benchmark model to quantify the impact of hiring discrimination on generating racially disparate dynamics within these cycles.
+Finally, we examine the welfare consequences of hiring discrimination during business cycles.
 5.1 Racial disparities over the business cycles
-Table 5 summarizes the business cycle disparities in the labor market between black and white workers.
-The top panel presents the dynamics of the racial unemployment rate gap, and the bottom panel presents the dynamics of the average wage.
+Table 6 presents the business cycle disparities in labor market outcomes between black and white workers.
+The top panel shows the dynamics of the racial unemployment rate gap, and the bottom panel shows the dynamics of the average wage.
 Cyclicality is measured by the correlation between the variable and real GDP.
-Volatility is calculated by the standard deviations of the variable relative to the standard deviation of real GDP.
+Volatility is calculated by the standard deviation of the variable relative to the standard deviation of real GDP.
 Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference.
-We then log the values and apply the HP filter with a smoothing parameter 1600 to extract cyclical components.
+We then log the values and apply the HP filter with a smoothing parameter of 1600 to extract cyclical components.
 We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions.
@@
 Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage.
-This is because of the composition change of workers over the business cycles.
+This is due to the composition change of workers over the business cycles.
 Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994).
-For the same reason, we report the cyclical properties of the average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index.
-The wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics.
-The wage volatility is calculated by the relative standard deviations of the year-to-year change in wage from the wage regression.
+For the same reason, we report the cyclical properties of average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index.
+Wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics.
+Wage volatility is calculated by the relative standard deviations of the year-to-year change in wages from the wage regression.
 [^12] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors.
@@
 We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics.
-Black workers’ average wage is more volatile (2.54) than white workers (2.08).
+Black workers’ average wage is more volatile (2.54) than that of white workers (2.08).
 5.2 Augmented model with aggregate shocks
 We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2.
-The TFP shocks follows an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$.
-The output of the matched firm-worker pairs depends on the realizations of both aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
+The TFP shocks follow an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$.
+The output of the matched firm-worker pairs depends on the realizations of aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
 With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers across employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$).
@@
 (2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015.
-All other parameters stay the same as in Table 1.
+All other parameters remain the same as in Table 1.
 This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
@@
 We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock.
-Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycle as a linear combination of the impulse response and the TFP realizations.
+Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycles as a linear combination of the impulse response and the TFP realizations.
 [^13] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system.
@@
 Table 7 compares data and model moments on unemployment and wages.
-Similar to standard search models, the stochastic dynamics of our model are under the influence of the Shimer Puzzle (Shimer, 2005).
-The augmented benchmark model generates similar cyclicality in the unemployment rate gap as data but much less volatility in the unemployment rate gap.
+Similar to standard search models, the stochastic dynamics of our model are subject to the Shimer Puzzle (Shimer, 2005).
+The augmented benchmark model generates a similar cyclicality in the unemployment rate gap as the data but with much less volatility in the unemployment rate gap.
 Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility.
-However, qualitatively, we have a higher wage cyclicality and volatility for black workers than for white workers.
+However, qualitatively, we observe a higher wage cyclicality and volatility for black workers than for white workers.
 5.3 Impact of hiring discrimination over business cycle
-We compare the benchmark dynamics with the counterfactual model without p firms to examine the role of discriminatory hiring over aggregate shocks.
+We compare the benchmark dynamics with the counterfactual model without $p$ firms to examine the role of discriminatory hiring in the context of aggregate fluctuations.
 [^14] As with Table 4, the counterfactual economy is a one-sector economy with only $np$ firms.
@@
 5.3.1 Impulse responses
-Figure 3 plots the response of black and white workers’ labor market, wealth, and consumption outcomes after a 1% unanticipated expansionary TFP shock.
-We compare the benchmark model response to the counterfactual model without $p$ firms.
-Overall, our models generate countercyclical unemployment rates and procyclical wage, wealth, and consumption movements.
+Figure 3 illustrates the labor market, wealth, and consumption responses of black and white workers following a 1% unexpected expansionary TFP shock.
+We evaluate the benchmark model’s reaction against the counterfactual model that excludes $p$ firms.
+In general, our models produce countercyclical unemployment rates alongside procyclical patterns in wages, wealth, and consumption.
 Figure 3: Impulse responses
@@
 The red solid line denotes the responses from the benchmark model, and the blue dashed line denotes the responses from the counterfactual model without $p$ firms.
-In the model without $p$ firms, black workers’ unemployment rate decreases less than that of the benchmark economy after the expansionary shock.
+In the model without $p$ firms, the unemployment rate of black workers decreases less than that of the benchmark economy after the expansionary shock.
 This implies that their unemployment rate is less volatile without $p$ firms.
-This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities to all workers.
-For white workers, the difference is much smaller between the benchmark and the model without $p$ firms.
+This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities for all workers.
+For white workers, the difference between the benchmark and the model without $p$ firms is much smaller.
 Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model.
-This is because the $p$ sector has a low job destruction shock.
+[^15] This is a common property in the literature on dual labor markets.
+A single-market model often reverts to its steady state faster than a dual-market model (e.g.
+Horvath and Yang, 2022).
+This is because the $p$ firms have a low job destruction shock.
 Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
-The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without discriminatory firms.
-This suggests more volatile procyclical wage and wealth responses for black workers than for white workers with the presence of $p$ firms.
+The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without prejudiced firms.
+This suggests more volatile procyclical wage and wealth responses for black workers than for white workers in the presence of $p$ firms.
 Lastly, we don’t observe a discerning difference in the consumption responses with and without $p$ firms.
@@
 The results are summarized in Table 8.
-The counterfactual equilibrium without $p$ firms eliminates the racial unemployment rate gap, resulting in the disappearance of its cyclical properties.
-Similarly, the racial differences in the cyclicality and volatility of average wages are largely reduced.
-As shown in Table 7, in the benchmark model, the black-white unemployment rate gap exhibits strong countercyclicality (-0.6731) and volatility (0.3009), both of which vanish in the absence of $p$ firms.
-For average wages, the cyclicality for black workers decreases from 1.0749 in the benchmark to 1.0131 without $p$ firms, while for white workers it increases slightly from 0.9984 to 1.0115.
-Similarly, the volatility for black workers falls from 1.0755 to 1.0139, while for white workers it increases from 0.9985 to 1.0116.
-This convergence demonstrates that hiring discrimination is a primary driver of the labor market’s cyclical disparities between black and white workers.
-These findings align with Cajner et al.
-(2017), who argue that demographic differences alone cannot fully account for observed racial disparities over the business cycle.
+In a counterfactual equilibrium without $p$ firms, the racial unemployment rate gap closes, eliminating its cyclical characteristics.
+Likewise, the disparities in the cyclicality and volatility of average wages between racial groups are significantly diminished.
+According to Table 8, the benchmark model reveals a pronounced countercyclicality (-0.6731) and volatility (0.3009) for the black-white unemployment rate gap, both of which disappear when $p$ firms are not included.
+For average wages, the cyclicality for black workers lowers from 1.0749 in the benchmark to 1.0131 without $p$ firms, whereas it increases slightly for white workers from 0.9984 to 1.0115.
+Conversely, the volatility for black workers drops from 1.0755 to 1.0139, while for white workers, it rises from 0.9985 to 1.0116.
+This convergence indicates that hiring discrimination significantly contributes to the cyclical disparities in the labor market between black and white workers.
+These results are consistent with Cajner et al.
+(2017), Ganong et al.
+(2020), and Ragusett (2022), which argue that demographic differences alone do not sufficiently explain the racial disparities observed throughout the business cycle.
 Table 8: Impact of hiring discrimination on business cycle racial gaps
@@
 Consumption ratio 0.9717 0.9695 0.1861 0.1659
-Note: The top panel compares the black-white unemployment rate, wage, consumption, and wealth gaps over the business cycles between the benchmark model and model without $p$ firms.
-We further examine the behaviors in the cyclical properties of the average black-to-white wealth ratio and the average black-to-white consumption ratio.
-Other than the cyclicality of the average wealth ratio, the cyclical properties of the racial gaps in all other dimensions are reduced after discriminatory hiring is removed.
+Note: The top panel compares the black-white unemployment rates, wages, consumption, and wealth gaps over the business cycles between the benchmark model and the model without $p$ firms.
 We further examine the cyclical properties of the average black-to-white wealth ratio and consumption ratio.
@@
 We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type.
-Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$.
+Low, mid, and high productivity correspond to the lowest, middle, and highest values of idiosyncratic productivity $s$.
 The wealth quintiles are based on benchmark steady-state wealth distribution.
-The economy experiences an average of 0.01% welfare decrease after removing prejudiced firms.
+The economy experiences an average welfare decrease of 0.01% after removing prejudiced firms.
 This is much less than the welfare change in the steady state in Table 5, primarily because workers are risk averse.
-All workers experience a decrease in unemployment rate volatility, and white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions.
+All workers experience a decrease in unemployment rate volatility, while white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions.
 On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%.
@@
 We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation.
-Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in steady state and business cycles.
+Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in the steady state and business cycles.
 Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum.
@@
   5. 5.
-    Unemployment insurance tax rate $\tau$
+[…]
   6. 6.
@@
   6. 6.
-    Unemployment insurance tax rate $\tau$ solves to balance the government budget.
+    Labor income tax rate $\tau$ solves to balance the government budget of the total unemployment insurance payout.
   7. 7.
@@
 Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks.
-Equalizing black workers’ bargaining power to white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes.
-This is because non-prejudiced firms retain less profit, hence posting fewer vacancies.
-On the net, it also reduces white worker’s job-finding rate and average wage rate.
+Equalizing black workers’ bargaining power with that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes.
+This is because non-prejudiced firms retain less profit, hence post fewer vacancies.
+On the net, it also reduces white workers’ job-finding rate and average wage rate.
 As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
@@
 In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers.
-The effect resembles assigning a higher bargaining power to black workers.
-This is because of the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012).
+The effect resembles assigning higher bargaining power to black workers.
+This is due to the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012).
 With a lower probability of losing their wealth, black workers can accumulate more personal wealth.
-Higher personal wealth gives black workers higher reservation value when bargaining with firms.
-Effectively, black workers can bargain for higher wage outcomes (1.66 compared to 1.60 in the benchmark model).
-Similar to directly assigning a higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent.
+Higher personal wealth gives black workers a higher reservation value when bargaining with firms.
+Effectively, black workers can negotiate for better wage outcomes (1.66 compared to 1.60 in the benchmark model).
+Similar to directly assigning higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent.
 The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output.
 It spills over to an increase in the average wage rate for all.
-Though it has a much less impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap.
+Though it has a much lesser impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap.
 Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).

### v3

@@
 Using a labor search-and-matching model with racially prejudiced and non-prejudiced firms, we show that labor market frictions sustain discriminatory practices as an equilibrium outcome.
-These practices account for 44% to 52% of the average wage gap and 16% of the median wealth gap.
-Discriminatory hiring also amplifies unemployment and wage volatility for black workers, increasing their labor market risks over business cycles.
-Eliminating prejudiced firms reduces these disparities and improves the welfare of black workers, though it slightly decreases overall economic welfare.
+These practices account for 57% of the racial unemployment gap, 48% of the average wage gap, and 16% of the median wealth gap.
+Discriminatory hiring also increases unemployment and wage volatility for black workers, increasing their labor market risks over the business cycle.
+Eliminating prejudiced firms reduces these disparities and improves the welfare of black workers as well as the overall economic welfare.
 JEL classification: D14, E21, J15, J64, J65
-Keywords: search-and-matching, heterogeneous agents, racial inequality, unemployment, wealth distribution, business cycles
+Keywords: search-and-matching, heterogeneous agents, hiring discrimination, racial inequality, labor market frictions, unemployment, wage gaps, wealth distribution, business cycles
 1 Introduction
-Discrimination is often considered inefficient or a response to information asymmetry.
-[^1] See Becker (1957) and Lang and Spitzer (2020) for discussions on taste-based and statistical discrimination theories.
-Market competition and better information are expected to eliminate discrimination over time.
+Market competition and increased access to information are expected to eliminate discrimination over time.
 Nonetheless, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities.
@@
 (2020); Derenoncourt and Montialoux (2021); Lippens et al.
-(2023).
-Resume studies consistently reveal racial discrimination even when controlling for candidate qualifications (e.g., Bertrand and Mullainathan, 2004; Bertrand and Duflo, 2017; Kline et al., 2022).
+(2023); Lang and Spitzer (2020).
+Resume studies consistently reveal racial discrimination even when controlling for candidate qualifications (e.g., Bertrand and Mullainathan, 2004; Bertrand and Duflo, 2017; Quillian et al., 2017; Kline et al., 2022).
 Gaps in labor income and wealth between black and white households have persisted long after the Civil Rights Movement (e.g., Cajner et al., 2017; Derenoncourt et al., 2023).
@@
 (2021).
-Yet, the role of discriminatory practices in shaping these disparities remains under-explored.
+Yet, the role of discriminatory practices in shaping these disparities remains underexplored.
 Developing a general equilibrium framework to explain the persistence of discrimination and its consequences is crucial for understanding the macroeconomic impacts of racial disparities and informing effective policymaking.
-This paper constructs a labor search-and-matching framework to capture the persistence of hiring discrimination based on race.
-We show that hiring discrimination accounts for 44% to 52% of the black-to-white average wage gap and 16% of the racial median wealth gap.
-Additionally, hiring discrimination increases both unemployment and wage volatility among black workers.
-It creates a countercyclical racial unemployment gap and diverges cyclical wage patterns between races.
-The model incorporates two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally.
-We include additional racial differences in the model as differences in bargaining power and the likelihood of experiencing wealth destruction.
-While all workers compete for jobs, white workers have additional opportunities with prejudiced firms.
+This paper constructs a search-and-matching model to examine the macroeconomic effects of persistent racial discrimination in hiring.
+We find that hiring discrimination explains roughly 57% of the black-white unemployment rate gap, 48% of the average wage gap, and 16% of the median wealth gap in the steady state.
+Moreover, discriminatory hiring amplifies the volatility of unemployment and wages among black workers over the business cycle.
+Eliminating hiring discrimination not only improves the welfare of black workers but also enhances overall economic welfare.
+The model features two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally.
+As a result, all workers compete for jobs at non-prejudiced firms, but white workers have the additional opportunity to match with prejudiced firms.
 Matched worker-firm pairs negotiate wages to maximize their joint surplus.
-Upon calibration, the model endogenously produces lower job-finding rates, higher unemployment rates, and lower bargained wages for black workers.
-Even without further financial frictions, the model generates significant racial disparities in wealth accumulation.
-The main message of this paper is that labor market friction can sustain hiring discrimination as an equilibrium outcome.
+Upon calibration, the model endogenously produces a lower job-finding rate, higher unemployment rate, and lower bargained wages for black workers.
+Even without additional financial frictions, the model generates substantial racial disparities in wealth accumulation.
+The main message of this paper is that labor market frictions can sustain hiring discrimination as an equilibrium outcome.
 Our calibration shows that prejudiced firms incur higher vacancy posting costs to discriminate against black workers.
-As we gradually increase these costs, prejudiced firms slowly reduce their vacancy postings and employment, while their profits adjust to satisfy the free-entry condition.
-Ultimately, we more than double the benchmark calibrated posting cost until the entire expected value of entering as a discriminatory firm is eliminated.
-This indicates that prejudiced firms maintain sufficient profit to absorb higher vacancy posting costs, even under competitive market conditions.
-Moreover, hiring discrimination perpetuates the equilibrium economy through firms’ vacancy postings.
-Prejudiced firms compete with non-prejudiced firms for unemployed white workers, driving up wage rates for white workers across the economy.
-This wage pressure reduces the expected profits of non-prejudiced firms, leading them to post fewer vacancies overall.
-The resulting decline in non-prejudiced job openings directly limits employment opportunities for black workers, contributing to lower wages, higher unemployment rates, and adverse wealth accumulation for them.
-In a counterfactual economy without prejudiced firms, the racial unemployment gap disappears, and the wage gap decreases significantly.
-Such a reduction in labor market disparity leads to asymmetric welfare changes.
-Black workers’ employment opportunities and wages increase because non-prejudiced firms expand their vacancy postings to fill positions left by prejudiced firms.
-As a result, black workers experience a welfare increase, with the most significant gains among high-productivity black workers.
-White workers, however, lose the exclusive job opportunities that prejudiced firms previously provided.
-Their average wage declines, and their unemployment rate rises.
-This leads to a moderate welfare decrease, especially among the lowest-productivity and lowest-wealth white workers.
-Due to asymmetrical job destruction shocks and vacancy posting costs, non-prejudiced firms cannot fully replace prejudiced firms.
-However, if we maintain the race-specific job destruction shocks, eliminating prejudiced hiring has a much smaller adverse impact on white workers.
-We further introduce aggregate shocks to the economy and show that hiring discrimination generates distinct business cycle dynamics for black and white workers.
-Our model successfully replicates the observed countercyclical and volatile racial unemployment rate gap, as well as the more procyclical and volatile average wage rate for black workers.
-Upon removing prejudiced firms, the business cycle patterns of the racial unemployment rate gap disappear.
-The volatility and cyclicality of average wages for black and white workers become much more similar.
-This provides theoretical evidence to Cajner et al.
-(2017) that firm-side discrimination contributes to the business cycle differences between black and white workers, and non-market demographic factors alone do not fully account.
-Lastly, we examine the associated business cycle welfare change between black and white workers when discriminatory hiring contributes to more volatile labor market outcomes for black workers.
-Not all workers benefit from a less volatile economy without prejudiced firms.
-Black workers gain an average of 0.02% in welfare, whereas white workers face a welfare loss of 0.01%.
-Different from the steady-state welfare values, among black workers, those with high productivity and high wealth gain the most.
-Among white workers, those with high productivity and high wealth lose the most, experiencing the largest welfare decreases within their group.
-This paper contributes to the ongoing discussion on the persistence of racial inequality by focusing on the disparate labor market conditions of white and black workers.
+Yet, they sustain positive profits by maintaining a lower turnover rate upon hiring white workers.
+Moreover, hiring discrimination affects the equilibrium economy through firms’ vacancy postings.
+Prejudiced firms compete with non-prejudiced firms for white workers, driving up their wage rates across the economy.
+This wage pressure reduces the expected profit of non-prejudiced firms, leading them to post fewer vacancies overall.
+The resulting decline in job openings at non-prejudiced firms directly limits employment opportunities for black workers, contributing to lower wages, higher unemployment, and adverse wealth accumulation.
+In a counterfactual economy without prejudiced firms, these racial disparities narrow substantially.
+Eliminating discriminatory hiring also yields significant welfare gains.
+Black workers experience a 12.9% increase in welfare as non-prejudiced firms expand their vacancy postings, which in turn improves their job opportunities and wages.
+Although white workers incur a modest 0.5% welfare decline due to the loss of exclusive job opportunities, the overall welfare of the economy rises by 1.5% because the gains for black workers more than offset the losses for white workers.
+We further incorporate aggregate shocks into our benchmark economy to examine how hiring discrimination shapes business cycle dynamics.
+Our model replicates key empirical patterns: a countercyclical and volatile racial unemployment gap, as well as a more procyclical and volatile wage response among black workers.
+In a counterfactual model without prejudiced firms, the volatility of unemployment rates and wages for black workers is substantially reduced relative to white workers.
+Moreover, hiring discrimination amplifies the volatility of racial disparities in consumption and wealth over the business cycle.
+These findings suggest that firm-side hiring discrimination plays a crucial role in driving the distinct business cycle dynamics of black and white workers, which cannot be fully explained by demographic factors alone (Cajner et al., 2017; Boulware and Kuttner, 2019, 2024).
+This paper contributes to the ongoing discussion on the persistence of racial inequality by focusing on the disparate labor market conditions of black and white workers.
 Numerous studies have documented racial discrimination in pay and employment opportunities that persist despite policy interventions (e.g., Coate and Loury, 1993; Black, 1995; Rosén, 1997; Bertrand and Mullainathan, 2004; Manduca, 2018).
@@
 (2020) suggest that limited employment opportunities for black workers contribute to the growing racial earnings gap.
-Our paper adds to this body of work by evaluating the impact of hiring discrimination on the cyclical aspects of racial disparities, examining how discriminatory practices exacerbate volatility and cyclicality in employment and wages for black workers.
+Recent research by Boulware and Kuttner (2019, 2024) further highlights that racial disparities intensify during downturns, driven by increased hiring discrimination and differential sensitivity of black employment to cyclical fluctuations, effects that cannot be fully attributed to black workers’ concentration in cyclically sensitive industries.
+Our paper adds to this body of work by evaluating the impact of hiring discrimination on the cyclical aspects of racial disparities, examining how discriminatory practices exacerbate the volatility and cyclicality of employment and wages for black workers.
 An emerging strand of literature explores racial disparities in wealth accumulation.
@@
 (2021) discuss how monetary policy can have disparate effects on workers of different racial backgrounds.
-Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over into long-term wealth accumulation disparities between black and white workers.
-Germane to our project, Nakajima (2021) develops a search-and-matching model to examine how monetary policies exacerbate racial differences in the labor market.
+Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over to long-term wealth accumulation disparities between black and white workers.
+Germane to our project, Nakajima (2021) develops a search-and-matching model to examine how monetary policy exacerbates racial differences in the labor market.
 Our model explicitly distinguishes between prejudiced and non-prejudiced firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy.
@@
 Section 4 examines the steady-state implications of racial discrimination.
-Section 5 provides the business cycle implications.
+Section 5 discusses the business cycle implications.
 Section 6 concludes the paper.
@@
 The model has a unit measure of workers, who are either black or white, $R\in\{bl,wh\}$.
-Two types of firms post vacancies to hire unemployed workers, subject to search and matching frictions.
-Prejudiced firms ($p$) actively discriminate against black workers and hire only white workers.
-Non-prejudiced firms ($np$) hire both black and white workers without discrimination.
-Prejudiced and non-prejudiced firms also experience type-specific job destruction shocks.
-Each firm posts one vacancy or hires one worker.
-Thus, firms and jobs are synonymous in our model.
-To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote the variables in the next period.
+Two types of firms post vacancies subject to search and matching frictions.
+Prejudiced firms ($p$) discriminate in hiring by effectively excluding black applicants, whereas non-prejudiced firms ($np$) hire workers without regard to race.
+A matched firm-worker pair corresponds to a filled job position, making firms and jobs synonymous within our model.
+Black workers thus receive offers exclusively from non-prejudiced firms, while white workers can match with both types of firms, increasing their employment opportunities.
+This assumption closely aligns with extensive empirical evidence on racial disparities in hiring.
+Audit studies by Bertrand and Mullainathan (2004) find that resumes with white-sounding names receive approximately 50% more callbacks than identical resumes with black-sounding names.
+Similarly, a meta-analysis by Quillian et al.
+(2017) confirms that white applicants consistently receive, on average, 36% more callbacks.
+Although discrimination occurs in other aspects such as wages, promotions, and separations, we specifically focus on hiring discrimination due to its documented significance, allowing us to clearly evaluate its distinct macroeconomic effects.
 Workers are either employed (with $p$ or $np$ firm) or unemployed.
@@
 The workers face idiosyncratic productivity shocks $s$, following an AR(1) process $\log(s^{\prime})=\rho_{s}\log(s)+\epsilon_{s}$, with $\epsilon_{s}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{s}^{2})$.
+[^3] To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote variables in the next period.
 Workers also differ in their asset holdings.
-They can save using these risk-free assets to partially insure themselves against the labor market risks.
+They can save using risk-free assets to partially insure themselves against the labor market risks.
 Following Mukoyama (2013), workers also receive race-specific extreme wealth shocks with probability $\epsilon_{R}\in\{\epsilon_{bl},\epsilon_{wh}\}$.
-Upon realization of the shock, the worker loses all their wealth.
+Upon realization of the shock, a worker loses all their wealth.
 Altogether, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$).
 The endogenous distribution of workers is summarized by $\mu(e,R,s,a)$.
-All endogenous aggregate variables depend on $\mu$.
-Individual workers may move around the distribution between periods; however, the overall distribution remains stationary.
-We omit the record of $\mu$ in the steady state and will reintroduce $\mu$ when augmenting the model with aggregate shocks in Section 5.
+In the steady state, individual workers may move around the distribution, but the overall distribution remains stationary.
+Thus, $\mu$ is not a state variable in the steady state model but becomes an aggregate state variable in the augmented model with aggregate shocks discussed in Section 5.
 2.2 Labor market search and matching
-The total number of unemployed workers, denoted as $u$, is the sum of unemployed black workers $(u_{bl})$ and unemployed white workers $(u_{wh})$.
+The total number of unemployed workers denoted as $u$, is the sum of unemployed black workers $(u_{bl})$ and unemployed white workers $(u_{wh})$.
 There are $v_{np}$ vacancies available in the non-prejudiced market, while the number of vacancies in the prejudiced market is $v_{p}$.
@@
 (2018), unemployed workers and vacant firms match through a constant returns to scale matching function
-$M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}},$ (1)
+$M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}}$ (1)
 with $\iota>0$.
 As documented by Den Haan et al.
-(2000), this functional form ensures that the matching probabilities lie within 0 and 1.
+(2000), this functional form ensures that the matching probabilities stay within 0 and 1.
 The probability for an unemployed worker to match with a vacant $np$ firm is $f(\theta_{np})=M(u,v_{np})/u=(1+\theta_{np}^{-\iota})^{-1/\iota}$, while the probability that a white unemployed worker matches with a vacant $p$ firm is $f(\theta_{p})=M(u_{wh},v_{p})/u_{wh}=(1+\theta_{p}^{-\iota})^{-1/\iota}$.
 Correspondingly, the probability of filling a vacant $np$ firm is $q(\theta_{np})=M(u,v_{np})/v_{np}=(1+\theta_{np}^{\iota})^{-1/\iota}$, while the probability of filling a vacant $p$ firm is $q(\theta_{p})=M(u_{wh},v_{p})/v_{p}=(1+\theta_{p}^{\iota})^{-1/\iota}$.
-In addition, non-prejudiced matches get separated with probability $\lambda_{np}$, while the prejudiced matches separate at the rate $\lambda_{p}$.
+In addition, non-prejudiced matches get destructed with probability $\lambda_{np}$, while the prejudiced matches separate at the rate $\lambda_{p}$.
 2.3 Unemployment Insurance
@@
 2.4 Workers
-The value function of an employed individual with race $R$, productivity $s$, asset $a$, and working with a $np$ firm is given by $W_{np}(R,s,a)$, while that of a white worker employed with a $p$ firm is given by $W_{p}(wh,s,a)$.
-Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced ($U^{I}_{np}(R,s,a)$) or a prejudiced ($U^{I}_{p}(wh,s,a)$) firm previously.
-An unemployed worker who is not eligible for unemployment benefits has $U^{N}(R,s,a)$ over their lifetime.
+The value function of an employed individual with race $R$, productivity $s$, asset $a$, and working with an $np$ firm is given by $W_{np}(R,s,a)$, while that of a white worker employed with a $p$ firm is given by $W_{p}(wh,s,a)$.
+Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced (with value function $U^{I}_{np}(R,s,a)$) or a prejudiced (with value function $U^{I}_{p}(wh,s,a)$) firm previously.
+An unemployed worker who is not eligible for unemployment benefits has a value of $U^{N}(R,s,a)$ over their lifetime.
 All the workers discount their future utility by $\beta$.
@@
 2.4.1 Employed with np firm
-Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in a $np$ firm.
+Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in an $np$ firm.
 The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their discounted lifetime utility.
-Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$.
-The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of extreme wealth shock $\epsilon_{R}$.
+Their income consists of period wage, $\omega_{np}$, net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$.
+The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of the extreme wealth shock $\epsilon_{R}$.
 $W_{np}(R,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})\hat{W}_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}\hat{W}_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (2)
@@
 $c+a^{\prime}$ $=(1-\tau)\omega_{np}(R,s,a)+(1+r)a+d$
-The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3.
-If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$.
-Otherwise, one may stay employed with probability $1-\lambda_{np}$.
-If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
+The worker receives a job destruction shock specific to $np$ firms and becomes unemployed with probability $\lambda_{np}$.
+Otherwise, they continue to remain employed with probability $1-\lambda_{np}$.
+If unemployed, the worker receives unemployment benefits and realizes a value of $U^{I}_{np}$.
+Thus, the future value of an employed worker conditional on the realization of wealth shock is given by Equation 3.
 $\hat{W}_{np}(R,s^{\prime},a^{\prime})=\lambda_{np}U^{I}_{np}(R,s^{\prime},a^{\prime})+(1-\lambda_{np})W_{np}(R,s^{\prime},a^{\prime})$ (3)
 2.4.2 Employed with p firm
-Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, and we only keep track of white workers since prejudiced firms do not hire black workers.
+Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, but we only consider white workers since prejudiced firms do not hire black workers.
 $W_{p}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{W}_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{W}_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (4)
@@
 A worker losing the job in the current period is eligible for unemployment benefits and earns value $U^{I}_{p}$ in the next period.
-For each $s^{\prime}$ and after realizing $\epsilon_{wh}$, an employed worker’s future value expands to Equation 5.
+Conditional on realizing the wealth shock, an employed worker’s future value expands to Equation 5.
 $\hat{W}_{p}(wh,s^{\prime},a^{\prime})=\lambda_{p}U^{I}_{p}(wh,s^{\prime},a^{\prime})+(1-\lambda_{p})W_{p}(wh,s^{\prime},a^{\prime})$ (5)
 2.4.3 Unemployed and eligible workers
-Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or a $np$ firm.
+Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or in an $np$ firm.
 In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
@@
 $c+a^{\prime}$ $=(1-\tau)b_{np}(bl,s,a)+(1+r)a+d$
-Their future value $\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})$ expands according to Equation 7:
+Their future value $\hat{U}^{I}_{np}(bl,s^{\prime},a^{\prime})$ expands to Equation 7.
 $\hat{U}^{I}_{np}(bl,s^{\prime},a^{\prime})=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))[P_{e}U^{I}_{np}(bl,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(bl,s^{\prime},a^{\prime})]$ (7)
-The unemployed worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$.
-Otherwise, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$.
-If one loses the benefit, their value function becomes $U^{N}(bl,s^{\prime},a^{\prime})$.
+The unemployed black worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$.
+Otherwise, they remain unemployed and continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$.
+If one loses the benefit, the worker receives a value of $U^{N}(bl,s^{\prime},a^{\prime})$.
 Unemployed white worker from $np$ firm
-Unemployed white workers are recruited by both $p$ and $np$ firms.
-Equation 8 describes their value function.
+Unemployed white workers are recruited by both $p$ and $np$ firms, and their value function is given by Equation 8.
 Like a black worker, a white worker encounters an extreme wealth shock with probability $\epsilon_{wh}$.
-They also have a budget constraint from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer.
+They also have a budget constraint depending on their post-tax unemployment insurance payout, current savings, and a lump-sum dividend transfer.
 $U^{I}_{np}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{U}^{I}_{np}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U}^{I}_{np}(wh,s^{\prime},0)\big{]}\Big{\}}$ (8)
@@
 $c+a^{\prime}$ $=(1-\tau)b_{np}(wh,s,a)+(1+r)a+d.$
-Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$.
-With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives an only job offer from a $p$ firm.
-With probability $f(\theta_{np})(1-f(\theta_{p}))$, they receive an only offer from a $np$ firm.
+Equation 9 describes the future value of an unemployed white worker conditional on the wealth shock.
+With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives a job offer only from a $p$ firm, while they receive an offer only from a $np$ firm with probability $f(\theta_{np})(1-f(\theta_{p}))$.
 With probability $f(\theta_{np})f(\theta_{p})$, they receive offers from both $np$ and $p$ firms simultaneously.
 If the worker receives both $p$ and $np$ offers, they will choose the job with a higher lifetime value.
-The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$).
-In this case, one may continue to receive unemployment benefits with probability $P_{e}$.
-Otherwise, one moves to a state of receiving no unemployment benefit with probability $(1-P_{e})$.
+The white worker will remain unemployed if they don’t receive any offer (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$).
+In this case, the worker continues to receive unemployment benefits with probability $P_{e}$, and lose their eligibility with probability $(1-P_{e})$.
 $\hat{U}^{I}_{np}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (9)
@@
 Equation 10 describes the problem faced by a white worker who last worked with a $p$ firm.
-The problem is identical to the previous case, except that they receive a payout of $b_{p}(wh,s,a)$ from the unemployment insurance.
+The problem is identical to the previous case, except that they receive an insurance payout of $b_{p}(wh,s,a)$.
 $U^{I}_{p}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})\hat{U}^{I}_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U}^{I}_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (10)
@@
 $c+a^{\prime}$ $=(1-\tau)b_{p}(wh,s,a)+(1+r)a+d$
-After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the unemployed worker’s future value is described in Equation 11. It is analogous to Equation 9.
+Analogous to the previous case, the unemployed worker’s future value conditional on the wealth shock is given by Equation 11.
 $\hat{U}^{I}_{p}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (11)
@@
 2.4.4 Unemployed and ineligible workers
-Ineligible workers have lost their unemployment insurance and are still searching for jobs.
-They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again.
-Therefore, their consumption relies solely on previous savings and lump-sum dividends, independent of their prior employment status.
+Ineligible workers are those who have lost their eligibility for receiving unemployment insurance and are still searching for jobs.
+Once ineligible, they no longer become eligible again.
+Therefore, their consumption relies solely on their savings and lump sum dividends, regardless of their prior employment status.
 Black worker
@@
 $c+a^{\prime}$ $=(1+r)a+d$
-And Equation 13 expands the future value of the black worker. One has a probability of $f(\theta_{np})$ of finding a job - only with a $np$ firm; otherwise, they remain unemployed and ineligible.
+Equation 13 expands the conditional future value of the black worker. The worker finds a job with an $np$ firm with probability $f(\theta_{np})$, and remains unemployed and ineligible otherwise.
 $\hat{U}^{N}(bl,s^{\prime},a^{\prime})$ $=f(\theta_{np})W_{np}(bl,s^{\prime},a^{\prime})+(1-f(\theta_{np}))U^{N}(bl,s^{\prime},a^{\prime})$ (13)
 White worker
-Equation 14 describes the problem of a white worker searching for jobs without unemployment insurance.
+Similarly, Equation 14 describes the problem of a white worker searching for jobs without unemployment insurance.
 $U^{N}(wh,s,a)$ $=\max_{c,a^{\prime}\geq 0}\Big{\{}u(c)+\beta\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{wh})\hat{U}^{N}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}\hat{U}^{N}(wh,s^{\prime},0)\Big{]}\Big{\}}$ (14)
@@
 $c+a^{\prime}$ $=(1+r)a+d$
-And Equation 15 expands the future value of the white worker in this scenario.
-It is nearly identical to Equation 11.
-With a probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$, the unemployed worker does not find any job and remains unemployed and ineligible.
+Equation 15 expands the conditional future value of the white worker in this scenario.
+It is almost identical to Equation 11 except for the fact that, in the event the unemployed worker doesn’t find any job (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$), the worker continues to remain ineligible for receiving unemployment benefits.
 $\hat{U}^{N}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (15)
@@
 2.5 Firms
-A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. These firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
+A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. All firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
 2.5.1 Vacant np firm
@@
 These firms match with an unemployed worker with probability $q(\theta_{np})$.
-Equation 16 describes the value of a vacant $np$ firm, $V_{np}$.
-$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}\Big{]}$ (16)
-$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}$
-$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))\Big{]}\Big{\}}da,$
-The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$.
-A $np$ firm can match with either a black or a white worker from the current unemployment pool.
+The value of a vacant $np$ firm, $V_{np}$ is given by Equation 16.
+$V_{np}$ $=-\kappa_{np}+\frac{q(\theta_{np})}{1+r}\int_{a}\Bigg{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{bl})J_{np}(bl,s^{\prime},g(bl,s,a))+\epsilon_{bl}J_{np}(bl,s^{\prime},0)\Big{]}\frac{\phi_{u}(bl,s,a)}{u}$ (16)
+$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{wh})\tilde{V}_{np}(wh,s^{\prime},g(wh,s,a))+\epsilon_{wh}\tilde{V}_{np}(wh,s^{\prime},0)\Big{]}\frac{\phi_{u}(wh,s,a)}{u}\Bigg{\}}da$
+where
+$\tilde{V}_{np}$ $(wh,s^{\prime},g(wh,s,a))=\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))\geq W_{p}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{np}(wh,s^{\prime},g(wh,s,a))\Big{)}$
+$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))<W_{p}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{np}(wh,s^{\prime},g(wh,s,a))\Big{)}(1-f(\theta_{p}))$
+with $g(R,s,a)$ denoting the worker’s optimal choice of asset $a^{\prime}$ for the next period.
+The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the expected value of production.
+An $np$ firm can match with either a black or a white worker from the current unemployment pool.
 $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers.
-Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ are the densities of the unemployed black and white workers over $s$ and $a$.
-These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$.
-Since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}})$, or if the worker did not receive a $p$ firm offer, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}(1-f(\theta_{p})))$.
+Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$.
+An unemployed black worker will accept the $np$ job offer once they get matched.
+On the other hand, since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, denoted by $\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))\geq W_{p}(wh,s^{\prime},g(wh,s,a))\}}$, or if the worker did not receive a $p$ firm offer, described by $\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))<W_{p}(wh,s^{\prime},g(wh,s,a))\}}(1-f(\theta_{p}))$.
 2.5.2 Vacant p firm
@@
 The probability of matching with an unemployed white worker is $q(\theta_{p})$.
-Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
-$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}$ (17)
-$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))\Big{]}da$
-Since the vacant $p$ firm only searches among white workers, the probability of matching with an unemployed white worker who has productivity $s$ and assets $a$ is represented by $\phi_{u}(wh,s,a)/u_{wh}$.
-Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than other options or if they do not receive a competing offer from an $np$ firm.
-We assume that there is free entry of firms, and hence both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
+Equation 17 describes the value of maintaining a $p$ vacancy, $V_{p}$.
+[…]
+where
+$\tilde{V}_{p}$ $(wh,s^{\prime},g(wh,s,a))=\mathbbm{1}_{\{W_{p}(wh,s^{\prime},g(wh,s,a))>W_{np}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{p}(wh,s^{\prime},g(wh,s,a))\Big{)}$
+$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},g(wh,s,a))\leq W_{np}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{p}(wh,s^{\prime},g(wh,s,a))\Big{)}(1-f(\theta_{np}))$
+Since the vacant $p$ firm searches only among white workers, the probability of matching with an unemployed white worker with productivity $s$ and asset $a$ is given by $\phi_{u}(wh,s,a)/u_{wh}$.
+Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than the other option or if they do not receive a competing offer from an $np$ firm.
+We assume that there is free entry of firms, and therefore both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
 2.5.3 Producing np firm
-The value of a producing $np$ firm is described by Equation 18.
-The firm chooses capital $k$ to maximize its lifetime value producing with the worker.
-The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$.
+The problem of a producing $np$ firm is described by Equation 18.
+The firm chooses capital $k$ to maximize its lifetime value, with the future value discounted using the interest rate and adjusted for the job destruction rate $\lambda_{np}$.
 Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm.
 The future value of the firm incorporates this probability.
-In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output minus production costs.
-The worker’s productivity, $s$, is a factor in the production function $sf(k)$.
-The matched firm produces $sf(k)$ units of output and incurs costs for capital rental and depreciation, represented by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
+In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output net of production costs.
+The matched firm produces $sf(k)$ units of output and incurs costs for renting capital and depreciation, denoted by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
 $J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\frac{1-\lambda^{np}}{1+r}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
@@
 $j(R,s,a)$ $=sf(k)-(r+\delta)k-\omega_{np}(R,s,a).$
-2.6 Producing p firm
+2.5.4 Producing p firm
 Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm.
-The firm rents capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
+The firm rents capital and discounts future profit accounting for the job destruction rate $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
 $J_{p}(wh,s,a)$ $=\max_{k}\Big{\{}j(wh,s,a)+\Big{(}\frac{1-\lambda^{p}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{wh})J_{p}(wh,s^{\prime},a^{\prime})+\epsilon_{wh}J_{p}(wh,s^{\prime},0)\big{]}\Big{\}}$ (19)
@@
 2.6 Wage bargaining
-In each period, matched worker-firm pairs bargain for wages over the matched surplus.
-A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$.
-A worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$.
-The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
+In each period, matched worker-firm pairs bargain for wages over the match surplus.
+A worker upon matching earns a surplus of $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ while the firm surplus is given by $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$.
+The bargaining power of the worker $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$.
+The resulting wage for workers employed at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
 $\omega_{np}(R,s,a)$ $=\operatorname*{arg\,max}_{\omega_{np}}\Big{(}W_{np}(R,s,a)-U^{I}_{np}(R,s,a)\Big{)}^{\xi_{R}}J_{np}(Ra,s,a)^{1-\xi_{R}}$ (20)
 $\omega_{p}(wh,s,a)$ $=\operatorname*{arg\,max}_{\omega_{p}}(W_{p}(wh,s,a)-U^{I}_{p}(wh,s,a))^{\xi_{wh}}J_{p}(wh,s,a)^{1-\xi_{wh}}$ (21)
-Finally, we define the stationary equilibrium in Appendix A.
+We define the stationary equilibrium in Appendix A.
 3 Calibration
 Each period in the model represents a quarter.
-We calibrate the benchmark model to match the relevant U.S.
-economy moments.
+We calibrate the benchmark model to match the relevant US economy moments.
 We have two sets of parameters.
-One group of parameters is chosen externally based on literature and empirical evidence without using model-generated data.
-The other parameters are calibrated internally by simulating our model to match a set of relevant data moments.
-We set the share of black workers to be 19.1% in the model, with the remaining being white workers.
-Table 1 shows the internally calibrated parameter values, their targeted moments, the externally chosen parameters, and their sources.
+One group of parameters is chosen externally based on the literature and empirical evidence without using model-generated data.
+The other set of parameters is calibrated internally by simulating our model to match a set of relevant data moments.
+Table 1 shows the values of the internal calibrated parameters, their targets, along with the externally chosen parameters and their sources.
+We set the percentage of black workers in the model at 14.9%, based on the Panel Study of Income Dynamics (PSID) data from 1996 to 2014.555Our sample follows Blundell et al.
+(2018).
+The share calculation follows Nakajima (2021).
 Table 1: Calibration and targeted statistics
@@
 $\beta$ 0.9943 discount factor K/Y 10.26 10.26
-$\iota$ 1.3012 matching elasticity job finding rate - black 0.4946 0.4946
-$\kappa_{p}$ 4.2622 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
-$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1.0 1.0
-$\lambda_{np}$ 0.0644 np sector job destruction shock job separation rate - black 0.0644 0.0644
-$\lambda_{p}$ 0.0268 p sector job destruction shock job separation rate - white 0.0380 0.0380
-$\xi_{bl}$ 0.1397 bargaining power - black mean wage ratio 0.75 0.75
-$\xi_{wh}$ 0.2110 bargaining power - white firm profit share 0.033 0.033
+$\iota$ 1.3452 matching elasticity job finding rate - black 0.4946 0.4946
+$\kappa_{p}$ 4.6314 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
+$\kappa_{\text{np}}$ 2.5534 np sector vacancy posting cost market tightness 1.0 1.0
+$\lambda_{np}$ 0.0644 np sector job destruction rate job separation rate - black 0.0644 0.0644
+$\lambda_{p}$ 0.0268 p sector job destruction rate job separation rate - white 0.0380 0.0380
+$\xi_{bl}$ 0.1371 bargaining power - black mean wage ratio 0.75 0.75
+$\xi_{wh}$ 0.1988 bargaining power - white firm profit share 0.033 0.033
 $\epsilon_{bl}$ 0.0179 extreme wealth shock - black zero wealth share - black 0.18 0.18
-$\epsilon_{wh}$ 0.0086 extreme wealth shock - white zero wealth share - white 0.07 0.07
+$\epsilon_{wh}$ 0.0088 extreme wealth shock - white zero wealth share - white 0.07 0.07
+$\chi$ 0.8403 maximum UI coverage fraction of median wage 0.48 0.48
+$Pe$ 0.5385 prob. of UI eligibility average weeks of eligibility 26 26
 Chosen externally
@@
 $h$ 0.4000 UI replacement rate Mitman and Rabinovich (2015)
-$\chi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
-$Pe$ 0.5385 probability of UI eligibility maximum weeks of eligibility
-Notes: This table reports the parameters, their values, and descriptions.
-The top panel presents the parameters chosen internally by minimizing the distance between model-generated moments and data.
-The last two columns of the top panel compare the targeted moments in the data and in the model simulations.
-The bottom panel reports the parameters chosen externally, their values, and description.
+Notes: This table presents the model parameters, their values, and descriptions.
+The top panel displays the parameters selected internally by minimizing the distance between the moments generated by the model and the moments derived from the data.
+The last two columns of the top panel compare the targeted moments in the data with those from the model simulations.
+The bottom panel provides the parameters chosen externally, their values, their descriptions and sources.
 3.1 Preferences
 We set the period utility function $u(c)$ to be $\log(c)$.
-The discount factor, $\beta$, is calibrated to match the quarterly capital-output ratio of 10.26, the value used by a number of studies including Den Haan et al.
+The discount factor, $\beta$, is calibrated to match the quarterly capital-output ratio of 10.26, a value used by a number of studies such as Den Haan et al.
 (2010) and Carroll et al.
@@
 3.2 Production
-The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$.
+The matched worker-firm pairs produce according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$.
 We choose $\alpha_{k}$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015, following Nakajima (2012).
 3.3 Productivity and wealth shocks
-We use hourly real wage from the Panel Study of Income Dynamics (PSID) to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process.
-Our estimation strategy closely follows Setty and Yedid-Levi (2021).
-We run a standard Mincer wage regression of log wages on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of men 25 years and older.
-We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data.
-The estimated regression is at a biennial frequency, given that the PSID data is available once every two years.
+We use hourly real wages from the PSID to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process.
+Our estimation strategy follows Setty and Yedid-Levi (2021).
+We run a Mincer regression of log wages on demographic controls and year and state fixed effects.
+The demographic controls include education, labor market experience, race, and marital status.
+We restrict the sample to men aged 25 and older.
+We then use the residuals obtained to estimate an AR(1) regression using the panel dimension of the PSID data.
+The estimated regression operates on a biennial frequency since the PSID data is released every two years.
 We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process.
 The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency.
-We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.555We follow Setty and Yedid-Levi (2021) to have only men in the sample.
-We also calculate the AR(1) processes, including men and women from PSID.
-The resulting $\rho_{s}$ is unchanged from the men-only sample (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
-Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth.
-The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers.
-They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth (Nakajima, 2021).
+We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, to be 0.1680.666We follow Setty and Yedid-Levi (2021) to have only men in the sample.
+We also estimate the AR(1) process to include both men and women from PSID.
+The resulting $\rho_{s}$ is unchanged (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
+Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people with zero wealth.
+The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0088 for white workers.
+They capture the empirical moments that around 18% of black workers and 7% of white workers have zero wealth (Nakajima, 2021).
 3.4 Unemployment insurance
-The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$.
-We choose the unemployment insurance replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others.
-Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model.
-The eligibility probability $P_{e}$ is chosen to be 0.5385 to generate an average unemployment benefit duration of 26 weeks, as in Mitman and Rabinovich (2015).
+The unemployment insurance system in our model is characterized by the replacement rate, $h$, the maximum insurance payout, $\chi$, and the probability of maintaining eligibility status, $P_{e}$.
+We choose $h$ to be 0.4 following Shimer (2005) and Mitman and Rabinovich (2015).
+In accordance with Setty and Yedid-Levi (2021), we calibrate $\chi$ to 0.8403, which amounts to 48% of the median wage in the model.
+The eligibility probability $P_{e}$ is set to 0.5385 to generate an average duration of unemployment benefits of 26 weeks, as in Mitman and Rabinovich (2015).
 3.5 Labor search and wage bargain
-The labor market turnover statistics are obtained from Cajner et al.
+The parameters governing the labor search are calibrated to match the labor market turnover statistics obtained from Cajner et al.
 (2017).
-The elasticity of the matching function, $\iota$, targets the job-finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al.
+The elasticity of the matching function, $\iota$, targets the job-finding rate of black workers.
+Our calibrated value of 1.3452 is close to 1.25, the value used by Den Haan et al.
 (2000) and Petrosky-Nadeau et al.
 (2018).
-We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively.
+We choose firm-type specific job destruction rates $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively.
 The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively.
-[^6] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers.
-On the other hand, white workers work in both $p$ and $np$ firms; hence, $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380).
-The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, of 1.0, following Wolcott (2021).
-[^7] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$.
-We find that the prejudiced firms pay a lot more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated to be 4.2622, compared to $\kappa_{np}$ taking a value of 2.735.
+[^7] All the employed black workers only work with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of black workers.
+White workers are employed in both $p$ and $np$ firms.
+Hence, $\lambda_{p}$ (0.0268) is lower than the aggregate separation rate of white workers (0.0380).
+The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers.
+The posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to match the aggregate labor market tightness $\theta$, following Wolcott (2021).
+[^8] Aggregate labor market tightness is defined as $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$.
+We find that prejudiced firms pay significantly more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated at 4.6314, compared to $\kappa_{np}$ at 2.5534.
 The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021).
-The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012).
-Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.2110.
+The bargaining power of the white worker, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012).
+Consistent with our expectations, black workers have a lower bargaining power than white workers, with $\xi_{bl}$ calibrated to be 0.1371 and $\xi_{wh}$ to be 0.1988.
 3.6 Assessing the model as a quantitative theory of racial disparity
 Our calibration successfully captures the racial gaps across three important dimensions: income, wealth, and employment outcomes.
-We now discuss these dimensions in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
+We now discuss the model validity in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
 (a) Labor income share
 (b) Wealth share
-Figure 1: Labor Income and Wealth Distribution
-Note: This figure compares the model-generated steady-state distributions of wealth and labor income with their empirical analogs.
-The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF).
-First, the model captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not target these distributions explicitly.
-As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile.
+Figure 1: Labor income and wealth distributions
+Note: This figure compares the steady-state distributions of wealth and labor income generated by the model with their empirical counterparts.
+The empirical labor income and wealth distributions are estimated using data from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF), respectively.
+The horizontal axes indicate quintiles, from the lowest (Q1) to the highest (Q5), while the vertical axes report the share of total income or wealth accounted for by each quintile.
+Blue bars represent the data, and red dotted bars show the model-simulated moments.
+First, the model successfully captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not explicitly target these distributions.
+As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile of the distribution.
 On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
 Unemp rate (White) 0.05 0.05
-Mean wealth ratio 0.23 0.28
+Mean wealth ratio 0.23 0.29
 Median wealth ratio 0.17 0.33
-Note: This table compares the steady state wealth and unemployment moments between black and white workers with the corresponding empirical data.
+Note: This table compares the empirical unemployment and wealth moments with the corresponding steady-state moments generated by the model.
 Second, the model reproduces the disparities in unemployment, labor income, and wealth between black and white workers.
-As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth.
-In addition to the targeted moments, Table 2 demonstrates that our model successfully generates the untargeted employment and wealth moments.
+Through calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth.
+In addition to the targeted moments in calibration, Table 2 shows that our model successfully generates the untargeted employment and wealth moments close to the data.
 Specifically, our model produces a 5% unemployment rate among white workers compared to 12% among black workers.
 Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to prejudiced firms.
-On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al.
-(2014).
-However, the model understates the median wealth gap between black and white workers.
-The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in Kaplan et al.
-(2014).
+On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.29, close to 0.23 in the data (Kaplan et al., 2014).
+However, the model underestimates the median wealth gap between black and white workers.
+The median black worker holds 33% of the wealth of the median white worker in our model compared to the empirical value of 17% (Kaplan et al., 2014).
 In summary, the model reflects the racial disparities present in labor income, wealth, and unemployment outcomes.
-Next, we will investigate how racial discrimination in hiring contributes to these disparities.
+Next, we investigate how racial discrimination in hiring contributes to these disparities.
 4 Steady state results
 In this section, we first establish how a search framework sustains hiring discrimination in equilibrium.
-Then, we examine the steady-state impact of racial discrimination by comparing the benchmark economy to an alternative economy without hiring discrimination.
+Then, we examine the steady-state impact of racial discrimination by comparing the benchmark economy with an alternative economy without hiring discrimination.
 Lastly, we explore the heterogeneous welfare implications associated with eliminating hiring discrimination.
 4.1 Sustaining racial discrimination
-The labor search-and-matching framework allows firms to earn a profit from the surplus generated by search frictions.
-This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium.
-It contrasts with a canonical frictionless neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957).
-Table 3 shows firm performances in the steady state.
-$p$ firms have a total profit of 0.04, while $np$ firms have 0.03.
-This difference arises because $p$ firms experience a lower job destruction shock (0.0268) than $np$ firms (0.0644).
-Although $p$ firms incur higher vacancy posting costs (4.2622) to hire only white workers, they retain workers longer (employment level at 0.54) at lower posting rates (0.02).
-Conversely, $np$ firms face lower vacancy posting costs (2.735), but their high job destruction shock leads to a lower employment level (0.40) and a more frequent need for vacancy postings (0.05).
+The presence of search-and-matching frictions allows firms to earn a profit from the surplus generated by search frictions.
+This feature permits prejudiced and non-prejudiced firms to coexist in the equilibrium.
+This is in contrast to a canonical frictionless neoclassical model in which prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers, as discussed in Becker (1957).
 Table 3: Firms in the steady state
@@
 $p$ firm profit 0.04
-np firm profit 0.03
+np firm profit 0.02
 p firm vacancy 0.02
 np firm vacancy 0.05
-p firm employment 0.54
-np firm employment 0.40
-Note: This table shows steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
-Next, we examine the steady-state interplay between prejudiced and non-prejudiced firms.
-At the benchmark equilibrium, $p$ firms post a positive number of vacancies with sufficient expected profits to cover the posting costs.
-We gradually increase penalties for them by raising $\kappa_{p}$ until their expected operating profit no longer justifies vacancy postings.
-Figure 2 displays the profit, vacancy, and employment dynamics between the $p$ and $np$ firms.
-As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment.
-Conversely, with fewer $p$ firms competing in the market, $np$ firms see a steady increase in profits, vacancy postings, and employment.
-It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks.
-As $p$ firms’ profits decline, the profits of $np$ firms increase at a slower rate.
-Additionally, as $np$ firms replace $p$ firms, $np$ firms recruit from both black and white workers, rather than only white workers displaced from $p$ firms.
-This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
-In the limit, as $\kappa_{p}$ approaches the total expected value of an operating $p$ firm, posting vacancies exclusively for white workers becomes entirely unprofitable.
-As a result, all $p$ firms exit the market.
-The value of $\kappa_{p}$ at which $p$ firms exit entirely is 217% of the benchmark value of 4.2622.
-Meanwhile, $np$ firms become the sole employers in the economy.
-The gray dotted lines in Figure 2 illustrate a steady state where we construct a counterfactual model that structurally removes $p$ firms from the benchmark model.
-Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of structurally eliminating $p$ firms.
-Figure 2: Penalizing discrimination
-Note: This figure displays the dynamics between $p$ and $np$ firms as $\kappa_{p}$ increases.
-The top panel plots the change in profit, the middle panel plots the change in vacancy posting, and the bottom panel plots the change in employment.
-The red dotted line depicts the dynamics of $p$ firms, the blue solid line depicts the dynamics of $np$ firms, and the gray dotted line depicts the limit where no $p$ firm exists.
-The horizontal axes record the values of $\kappa_{p}$ from the benchmark $\kappa_{p}$ at 4.2622 to 217% of the benchmark value, at 9.2444.
-The vertical axes record the model-generated value of each variable.
+p firm employment 0.57
+np firm employment 0.37
+Note: This table presents steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
+Table 3 shows the performance of firms in the steady state.
+Prejudiced firms earn a total profit of 0.04, while the corresponding profit of non-prejudiced firms is 0.02.
+This difference arises because $p$ firms experience a lower job destruction shock (0.0268) compared to $np$ firms (0.0644).
+Even though $p$ firms incur higher vacancy posting costs (4.6314) to hire only white workers, they retain their workers for a longer period, reducing the number of vacancy postings.
+In contrast, $np$ firms face lower vacancy posting costs (2.5534), but their higher job destruction rate leads to a more frequent need for vacancy postings (0.05).
+In the steady state, $p$ firms sustain an employment level (0.57) higher than $np$ firms (0.37).
 4.2 Impact of hiring discrimination on racial inequality
-In this subsection, we compare the benchmark economy to those without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring.
-Table 4 presents the comparisons.
-All firms in the counterfactual models provide equal hiring opportunities to black and white workers.
-Column 2 removes the $p$ firms from the benchmark model, which leaves all workers with equal job-finding rates with the $np$ firms and equal job destruction shocks at the benchmark $\lambda_{np}$.
-In Column 3, we further adjust the job destruction shocks to be race-specific, with $\lambda_{wh}$ and $\lambda_{bl}$ designating the shocks for white and black workers, respectively.
-[^8] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates.
-This counterfactual economy only has one type of firm.
-We assign white workers a job destruction shock, $\lambda_{wh}$, of 0.0380 and black workers a destruction shock, $\lambda_{bl}$, of 0.0644 to maintain black and white workers’ job separation rates as in the calibration.
-Reassigning the job destruction shocks further isolates the impact of job separation.
+In this subsection, we analyze the distributional and aggregate effects of discriminatory hiring by comparing the benchmark economy to a counterfactual economy without prejudiced firms.
+The counterfactual economy contains only non-prejudiced firms, where both black and white workers retain their race-specific job separation rates from the benchmark calibration.
+By holding these separation rates constant, we isolate how discrimination in hiring - rather than differences in job separation - shapes labor market outcomes.
+[^9] The benchmark economy calibrates prejudiced and non-prejudiced firms’ job destruction shocks ($\lambda_{p}$ and $\lambda_{np}$) to match observed racial disparities in job separation rates.
+In the single-firm-type counterfactual, we preserve benchmark separation rates by assigning white workers a $\lambda_{wh}$ of 0.0380 and black workers a $\lambda_{bl}$ of 0.0644.
+Table 4 presents the comparison.
 Table 4: Impact of hiring discrimination
-Moments Benchmark Eliminate $p$ firms Race specific $\lambda$
+Moments Benchmark No $p$ firms
 Households
-job separation rate - black 0.06 0.06 0.06
-job separation rate - white 0.04 0.06 0.04
-job finding rate - black 0.50 0.59 0.62
-job finding rate - white 0.66 0.59 0.62
-unemp rate - black 0.12 0.10 0.09
-unemp rate - white 0.05 0.10 0.06
-mean wage - black 1.60 1.96 1.82
-mean wage - white 2.13 2.22 2.11
-mean wage ratio 0.75 0.88 0.86
-mean wealth ratio 0.28 0.33 0.33
-median wealth ratio 0.33 0.44 0.44
+job separation rate - black 0.06 0.06
+job separation rate - white 0.04 0.04
+job finding rate - black 0.49 0.64
+job finding rate - white 0.66 0.64
+unemp rate - black 0.12 0.09
+unemp rate - white 0.05 0.06
+mean wage - black 1.59 1.85
+mean wage - white 2.12 2.11
+mean wage ratio 0.75 0.87
+mean wealth ratio 0.29 0.34
+median wealth ratio 0.33 0.44
 Firms
-$p$ firm profit 0.04 - -
-np firm profit 0.03 0.06 0.06
-p firm vacancy 0.02 - -
-np firm vacancy 0.05 0.10 0.07
-p firm employment 0.54 - -
-np firm employment 0.40 0.90 0.94
+$p$ firm profit 0.04 -
+np firm profit 0.02 0.06
+$p$ firm vacancy 0.02 -
+np firm vacancy 0.05 0.07
+$p$ firm employment 0.57 -
+np firm employment 0.37 0.94
 Labor Market
-$p$ market tightness 0.40 - -
-$np$ market tightness 0.73 1.02 1.12
+$p$ market tightness 0.39 -
+$np$ market tightness 0.71 1.17
 Aggregate Outcomes
-Y 3.06 2.95 3.06
-K/Y 10.26 10.22 10.25
-average wage 2.04 1.96 2.05
-unemp rate 0.07 0.10 0.06
+Y 3.07 3.08
+K/Y 10.26 10.27
+average wage 2.04 2.07
+unemp rate 0.06 0.06
 Note: This table compares the benchmark steady state with a counterfactual steady state without $p$ firms.
-In the counterfactual model, all firms provide equal hiring opportunities to black and white workers and keep race-specific job destruction shock.
+In the counterfactual model, all firms provide equal hiring opportunities to black and white workers.
 All other parameters stay the same as benchmark calibration.
-The first panel illustrates the changes in racial employment, wage, and wealth disparities after removing prejudiced firms.
-The first two rows describe the job separation rates.
-By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the differences as the benchmark calibration.
-The next two rows describe the job-finding rates.
-Since there are no $p$ firms in the economy in Columns 2 and 3, the job-finding rates are equal for black and white workers.
-White workers have only one sector for employment; hence, their job-finding rate decreases.
-Black workers have more job opportunities as the $np$ sector posts more vacancies; thus, their job-finding rate increases.
-Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2.
-Altogether, removing $p$ firms reduces the racial gap in the unemployment rate.
-In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes.
-In Column 3, the unemployment rate gap closes from 7% to 3% due to the remaining racial difference in $\lambda$.
-Turning to the wage impact, black workers can bargain for higher wage rates as they receive more job opportunities when we remove discriminatory hiring.
-White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rates.
-As a result, the average black-to-white wage ratio increases from 75% in the benchmark to 88% in Column 2 and to 86% in Column 3.
-We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
-Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44.
-The results remain stable between Columns 2 and 3.
-We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
-The remaining gaps in wage and wealth come from the bargaining power and extreme wealth shock disparities between black and white workers.
-[^9] Though not the primary purpose of this paper, we present the results after removing racial differences in bargaining power and extreme wealth shock in Appendix B.
-The second panel displays the changes in firm attributes.
-[^10] We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock.
-Without prejudiced firms, $np$ firms’ profits double, and vacancy postings and employment levels increase.
-These changes result in the aforementioned advantages for black workers in the counterfactual economies.
-In Column 3, the lower job destruction shock for white workers further drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
-The third panel of Table 4 displays the labor market conditions.
-As $p$ firms exit, a greater number of $np$ firms enter the market endogenously.
-The labor market tightness for the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms, and to 1.12 when we further split $\lambda$ by race.
-The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
-The last panel of Table 4 describes the aggregate outcomes.
-With more $np$ firms entering, the aggregate output remains relatively unchanged in Column 3.
-The capital-output ratio slightly decreases to 10.25.
-The economy-wide average wage increases to 2.05 from the benchmark of 2.04, and the aggregate unemployment rate decreases to 6%.
-However, Column 2 experiences a more significant reduction in output and average wage, along with a higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs instead of using resources for production.
-Our results show that the presence of prejudiced firms can explain sizable fractions of income and wealth inequality.
-Penalizing prejudiced firms primarily improves employment opportunities and wage outcomes for black workers; it also worsens the situation for white workers.
-Removing $p$ firms decreases job-finding rates and increases unemployment for white workers, lowering their bargaining positions and affecting their income and wealth.
-To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by black and white workers.
+The first panel shows how racial disparities in household outcomes change when prejudiced firms are removed.
+While the racial differences in job separation rates remain unchanged at benchmark levels, eliminating discriminatory hiring practices leads to equalized job-finding rates for black and white workers.
+White workers experience a slight decrease in their job-finding rate, dropping from 0.66 to 0.64, due to losing access to $p$ firms.
+Meanwhile, black workers’ job-finding rate rises significantly, from 0.49 to 0.64, in the counterfactual equilibrium.
+The removal of $p$ firms leads to endogenous entry of $np$ firms, which increases the number of vacancies and thus raises the job-finding rate for black workers.
+Overall, this removal in prejudiced firms narrows the racial unemployment gap from 7 percentage points to 3 percentage points, showing that discriminatory hiring explains approximately 57% of the unemployment gap between black and white workers.
+Turning to the impact on wages, removing hiring discrimination boosts outcomes for black workers by improving their outside options.
+This enables black workers to bargain for higher wages upon a successful match.
+However, white workers no longer have access to jobs from the exclusive $p$ firms.
+This lowers their outside option during bargaining, slightly reducing their wage rates.
+As a result, the average wage ratio between black and white workers increases from 0.75 to 0.87.
+We can conclude that hiring discrimination accounts for roughly 48% of the wage gap.
+The change in hiring practices also narrows the wealth gap: the mean wealth ratio rises from 0.29 to 0.33, and the median wealth ratio increases from 0.34 to 0.44.
+These shifts imply that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
+The remaining differences in wages and wealth can be accounted for by differences in bargaining power and extreme wealth shocks between black and white workers.
+[^10] Although not the primary focus of this paper, additional results that remove racial differences in bargaining power and extreme wealth shocks are presented in Appendix B.
+The second panel examines firm dynamics when discriminatory firms are removed.
+[^11] We refer to the one-sector economy without discriminatory hiring as $np$ firms, although their job destruction rate $\lambda$ differs from that of the benchmark $np$ firms.
+Without prejudiced firms, $np$ firms see their profits triple, which encourages additional $np$ firms to enter the market and post vacancies.
+Consequently, $np$ firms fully replace $p$ firms in the counterfactual equilibrium.
+As shown in Table 4, the total number of vacancies and overall employment remain unchanged compared to the benchmark.
+Still, eliminating prejudiced firms helps narrow the racial gap in labor market outcomes, as described earlier.
+The third panel of Table 4 shows how labor market conditions adjust when $p$ firms exit.
+As these firms leave, additional $np$ firms enter endogenously, raising the labor market tightness in the $np$ market from 0.71 in the benchmark to 1.17 in the counterfactual equilibrium.
+Consequently, the overall market tightness exceeds the benchmark aggregate level of 1.
+The last panel of Table 4 presents the aggregate outcomes.
+Despite the increased entry of $np$ firms, aggregate output rises only marginally from 3.07 to 3.08.
+The capital-output ratio increases from 10.26 to 10.27, and the aggregate wage grows from 2.04 to 2.07, while the aggregate unemployment rate remains steady at 6%.
+Our results suggest that prejudiced firms account for a substantial share of employment, income, and wealth inequality.
+Eliminating $p$ firms improves employment and wage outcomes for black workers, while white workers experience a slight decline in their job-finding rate and an increase in unemployment.
+This reduction in job prospects weakens white workers’ bargaining power and adversely affects their income and wealth.
+In the next section, we quantify these differential effects by calculating the welfare changes experienced by black and white workers.
 4.3 Welfare analysis
 We follow Krusell et al.
-(2010) in measuring the welfare change for black and white workers by calculating the change in average consumption equivalence after removing discriminatory hiring.
-[^11] Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$.
-Under an alternate economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state.
-We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$.
-With $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$.
-We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change.
-We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms.
-Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift.
-We present the average welfare changes for black and white workers by productivity types and wealth quintiles in Table 5.
-Overall, the economy experiences a 5.83% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race.
-Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%.
-This is because of the rise in their job opportunities and higher wage rates.
-On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%.
-This is because of the decrease in their wage rates and the increase in the unemployment rate.
-Mid-panels in Table 5 present the heterogeneous welfare change for black and white workers by productivity types.
-Though the welfare reduction from eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms.
-In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race.
-Furthermore, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
+(2010) in measuring the welfare effects of discriminatory hiring on black and white workers.
+To do so, we compute the change in average consumption equivalence after removing $p$ firms.
+[^12] See Appendix C for details on the welfare calculation.
+Table 5 presents the overall welfare change, as well as the breakdown by productivity type and wealth quintile for both groups.
+Eliminating hiring discrimination raises aggregate welfare by 1.5%.
+Black workers experience a 12.9% increase in their average welfare, mainly due to improved job opportunities and higher wage rates.
+White workers, on the other hand, experience a modest 0.48% decline, driven by a slight drop in wages and a higher unemployment rate after losing access to the exclusive $p$ firms.
+Although white workers make up about 85% of the population, the substantial gains for black workers more than offset the small loss among white workers, resulting in a net rise in overall welfare.
 Table 5: Heterogeneous welfare change
-Average welfare gain (%) Eliminate $p$ firms Race specific $\lambda$
-Overall -5.83 1.17
-Black White Black White
-6.89 -8.83 10.94 -1.14
+Average welfare gain (%) No $p$ firms
+Aggregate 1.51
+Black White
+Overall 12.89 -0.48
 by productivity
-Low 6.30 -9.03 10.31 -1.07
-Mid 6.92 -8.83 10.98 -1.14
-High 7.41 -8.61 11.49 -1.19
+Low 12.11 -0.47
+Mid 12.94 -0.48
+High 13.57 -0.47
 by wealth
-Low 20% 6.85 -9.43 10.98 -1.13
-40-60% 6.95 -8.95 10.94 -1.17
-Top 20% 6.88 -8.14 10.72 -1.09
-Note: This table compares the average consumption equivalence change from the benchmark steady state to the equilibrium where we eliminate $p$ firms (Columns 1 and 2) and further adjust the job destruction shock to be race-specific (Columns 3 and 4).
+Low 20% 12.91 -0.45
+40-60% 12.92 -0.49
+Top 20% 12.65 -0.48
+Note: This table compares the average consumption equivalence change from the benchmark steady state to an equilibrium with no $p$ firms.
 Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$.
 The wealth quintiles are based on benchmark steady-state wealth distribution.
-The bottom panels of Table 5 present the welfare change by wealth quintiles for black and white workers.
-Removing prejudiced firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race.
-The most considerable reduction in welfare (9.43%) occurs for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
+The middle panel in Table 5 displays the heterogeneous welfare changes for black and white workers by productivity.
+For white workers, the welfare decline from eliminating prejudiced firms is similar across productivity levels, with a slightly larger loss for those in the middle productivity group.
+In contrast, among black workers, the highest productivity group experiences the greatest gain (13.57%), followed by the middle and low productivity groups.
+The bottom panel in Table 5 breaks down the welfare changes by wealth quintile.
+Removing prejudiced firms benefits black workers most in the middle quintile, where welfare rises by 12.92%, while the largest welfare decline for white workers also occurs in the middle quintile.
 5 Business cycle dynamics
-In this section, we begin by documenting the racial disparities over business cycles.
-Next, we incorporate aggregate uncertainty into our benchmark model to quantify the impact of hiring discrimination on generating racially disparate dynamics within these cycles.
-Finally, we examine the welfare consequences of hiring discrimination during business cycles.
+In this section, we first document the racial disparities over the business cycle.
+We then introduce aggregate uncertainty into our benchmark model to assess how hiring discrimination shapes racially disparate dynamics during economic fluctuations.
 5.1 Racial disparities over the business cycles
-Table 6 presents the business cycle disparities in labor market outcomes between black and white workers.
-The top panel shows the dynamics of the racial unemployment rate gap, and the bottom panel shows the dynamics of the average wage.
-Cyclicality is measured by the correlation between the variable and real GDP.
-Volatility is calculated by the standard deviation of the variable relative to the standard deviation of real GDP.
-Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference.
-We then log the values and apply the HP filter with a smoothing parameter of 1600 to extract cyclical components.
-We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions.
-It has a relative volatility of 9.71.
+Table 6 displays the labor market outcome disparities between black and white workers over the business cycle.
+The top panel presents the dynamics of unemployment rates for both groups, while the bottom panel shows the evolution of average wages after adjusting for composition bias.
+In this table, volatility is defined as the standard deviation of a variable relative to that of real GDP, and cyclicality is measured as the correlation between the variable and real GDP.
+Unemployment rate: We use quarterly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the business cycle dynamics of unemployment rates.
+The statistics are computed using log deviations from an HP-filtered trend with a smoothing parameter of 1600.
+We find that both black and white unemployment rates are countercyclical, although the black unemployment rate is slightly less cyclical and less volatile than the white rate.
+These findings align with Nakajima (2021).
+In addition, the black-white unemployment rate gap is itself countercyclical: it widens during recessions and narrows during expansions, indicating that black workers see a larger unemployment increase in recessions but a faster decrease in expansions.
+This result is consistent with the countercyclical gap documented by Cajner et al.
+(2017).
 Table 6: Business cycle statistics, US data, 1996-2014
-Cyclicality Volatility
+Volatility Cyclicality
 Unemployment rate
-Black-White gap -0.6255 9.7100
+Black 9.51 -0.85
+White 11.02 -0.92
+Black-White gap 9.46 -0.63
 Average wage
-Black 2.6041 2.5352
-White 1.2729 2.0848
+Black 2.54 2.60
+White 2.08 1.27
 Note: This table provides business cycle statistics for the US.
-The unemployment rate was obtained from the Current Population Survey (CPS) from 1996 to 2014.
-Average wage statistics are constructed from the Panel Study of Income Dynamics (PSID) of the same period.
-All variables are logged and HP filtered with a smoothing parameter of 1600.
-Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage.
-This is due to the composition change of workers over the business cycles.
-Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994).
-For the same reason, we report the cyclical properties of average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index.
-Wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics.
-Wage volatility is calculated by the relative standard deviations of the year-to-year change in wages from the wage regression.
-[^12] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors.
-Specifically, we follow Bils (1985), Solon et al.
-(1994), and Devereux (2001) and run the regression: $\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it}$ (22) where $Y_{t}$ refers to real GDP.
-Demographic variables $x_{it}$ include gender, education, labor market experience, and marital status.
-We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$.
-We measure cyclicality using coefficient $\beta$ from the regression.
-The volatility is the standard deviation of the year fixed effects, $\nu_{t}$.
-We then adjust the estimation to quarterly frequency.
-We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics.
-Black workers’ average wage is more volatile (2.54) than that of white workers (2.08).
+The unemployment rate is obtained from the Current Population Survey (CPS) for 1996 to 2014.
+Average wage statistics are constructed from the Panel Study of Income Dynamics (PSID) for the same period.
+Unemployment rate statistics are calculated using HP filtered series with a smoothing parameter of 1600, while wage statistics are obtained using wage growth rates after controlling for demographics.
+Wages: We use individual-level data from the PSID (1996–2014) to estimate the cyclical properties of black and white wages.
+Previous research (Stockman, 1983; Bils, 1985; Solon et al., 1994) shows that using aggregate data can be problematic due to changes in worker composition over the business cycle.
+Specifically, aggregate data tends to weight lower-skilled (and lower-wage) workers more heavily during expansions than recessions, causing countercyclical bias.
+To address this issue, we rely on the longitudinal structure of the PSID data to keep the composition of workers fixed over time.
+Following Bils (1985), Solon et al.
+(1994), and Devereux (2001), we estimate wage cyclicality by regressing wage growth on real GDP growth, controlling for observable characteristics.
+We measure wage volatility as the standard deviation of wage growth after controlling for demographics.
+[^13] Appendix D provides details on measuring the cyclicality and volatility of real wages.
+Our findings indicate that both black and white wages are procyclical, with black workers’ wages being twice as cyclical as those of white workers.
+Black workers’ wages also exhibit greater volatility than white workers’ wages.
 5.2 Augmented model with aggregate shocks
-We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2.
+We introduce shocks to aggregate total factor productivity (TFP) $z$ in our benchmark steady-state model described in Section 2.
 The TFP shocks follow an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$.
-The output of the matched firm-worker pairs depends on the realizations of aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
-With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers across employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$).
-The next period’s aggregate distribution $\mu^{\prime}$ is determined by $(z,\mu)$, and the law of motion is given by $\mu^{\prime}=\Gamma(z,\mu)$.
+The output of the matched worker-firm pairs depends on the realizations of both aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
+With the introduction of aggregate shocks, the state space expands to include aggregate states $(z,\mu)$, where $\mu$ is the distribution of workers over their employment status ($e$), race ($R$), idiosyncratic productivity ($s$), and asset ($a$).
+The aggregate distribution for the next period $\mu^{\prime}$ is determined by $(z,\mu)$, and the endogenous law of motion given by $\mu^{\prime}=\Gamma(z,\mu)$ is determined in the equilibrium.
 We follow Cooley et al.
 (1995) and Boppart et al.
-(2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015.
-All other parameters remain the same as in Table 1.
-This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
-We obtain the stochastic equilibrium of our model by using the sequence space method of Boppart et al.
+(2018) in choosing the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015.
+All other parameters remain the same as our benchmark calibration given in Table 1.
+We obtain the stochastic equilibrium of our model using the sequence space method proposed by Boppart et al.
 (2018).
@@
 [^14] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system.
-In Appendix C, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
+In Appendix E, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
 Table 7: Business cycle model vs data
-Cyclicality Volatility
+Volatility Cyclicality
 Data Model Data Model
 Unemployment rate
-Black-White gap -0.6255 -0.6731 9.7100 0.3009
+Black 9.51 0.26 -0.85 -0.69
+White 11.02 0.22 -0.92 -0.71
+Black-White gap 9.46 0.30 -0.63 -0.68
 Average wage
-Black 2.6041 1.0749 2.5352 1.0755
-White 1.2729 0.9984 2.0848 0.9985
-Note: This table compares the data moments and the benchmark generated moments on the black-white unemployment and wage gaps over the business cycles.
-Table 7 compares data and model moments on unemployment and wages.
-Similar to standard search models, the stochastic dynamics of our model are subject to the Shimer Puzzle (Shimer, 2005).
-The augmented benchmark model generates a similar cyclicality in the unemployment rate gap as the data but with much less volatility in the unemployment rate gap.
-Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility.
-However, qualitatively, we observe a higher wage cyclicality and volatility for black workers than for white workers.
+Black 2.54 1.08 2.60 1.08
+White 2.08 1.00 1.27 1.00
+Note: This table compares the data moments with the moments generated by the benchmark model on unemployment and wages.
+Table 7 compares business cycle statistics for unemployment and wages between the data and our model.
+Overall, the model replicates the key patterns observed in the data.
+It generates countercyclical unemployment rates, with white unemployment being slightly more cyclical than black unemployment, and it accurately reproduces the countercyclical gap between the two groups.
+In terms of wages, the model produces procyclical patterns, with black wages showing higher cyclicality and volatility than white wages.
+However, as is common with standard search models, our model falls short in generating realistic volatility levels for both black and white unemployment rates - a limitation discussed in Shimer (2005).
 5.3 Impact of hiring discrimination over business cycle
-We compare the benchmark dynamics with the counterfactual model without $p$ firms to examine the role of discriminatory hiring in the context of aggregate fluctuations.
-[^14] As with Table 4, the counterfactual economy is a one-sector economy with only $np$ firms.
-We do not further adjust race-specific destruction shocks.
+We compare the benchmark dynamics with a counterfactual model that excludes prejudiced firms to examine how discriminatory hiring impacts the economy under aggregate fluctuations.
+In this one-sector counterfactual, only non-prejudiced firms operate, and workers face race-specific job separation rates as with Table 4.
 5.3.1 Impulse responses
-Figure 3 illustrates the labor market, wealth, and consumption responses of black and white workers following a 1% unexpected expansionary TFP shock.
-We evaluate the benchmark model’s reaction against the counterfactual model that excludes $p$ firms.
-In general, our models produce countercyclical unemployment rates alongside procyclical patterns in wages, wealth, and consumption.
+Figure 2 shows how labor market outcomes, wealth, and consumption for both black and white workers respond to an unexpected 1% expansionary TFP shock.
+We compare the benchmark model’s reaction with that of a counterfactual model that excludes $p$ firms.
+In both models, unemployment rates move countercyclically while wages, wealth, and consumption follow procyclical patterns.
+However, the responses in the counterfactual economy are generally less pronounced.
 Figure 2: Impulse responses
-Note: This figure plots the impulse response functions of unemployment rates, black-white average wage and wealth ratios, and aggregate consumption between black and white workers to a one percent increase in aggregate TFP at date zero.
+Note: This figure plots the impulse response functions of unemployment rates, black-white average wage and wealth ratios, and aggregate consumption of black and white workers to a one percent increase in aggregate TFP at date zero.
 The red solid line denotes the responses from the benchmark model, and the blue dashed line denotes the responses from the counterfactual model without $p$ firms.
-In the model without $p$ firms, the unemployment rate of black workers decreases less than that of the benchmark economy after the expansionary shock.
-This implies that their unemployment rate is less volatile without $p$ firms.
-This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities for all workers.
-For white workers, the difference between the benchmark and the model without $p$ firms is much smaller.
-Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model.
-[^15] This is a common property in the literature on dual labor markets.
-A single-market model often reverts to its steady state faster than a dual-market model (e.g.
-Horvath and Yang, 2022).
-This is because the $p$ firms have a low job destruction shock.
-Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
-The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without prejudiced firms.
+In the model without $p$ firms, the unemployment rate for black workers falls less after an expansionary shock than in the benchmark, indicating that black unemployment is less volatile when discriminatory hiring is removed.
+In the benchmark economy, $np$ firms compete with $p$ firms for hires, which leads them to post more vacancies and create additional opportunities for workers.
+Although white workers follow a similar pattern, the difference between the benchmark and counterfactual responses is much smaller for them.
+Notably, the counterfactual model shows a much faster recovery of unemployment rates than the benchmark model.
+[^15] This is a common feature in dual labor market models, where a single-market setup typically reverts to its steady state faster than a dual-market framework (e.g., Horvath and Yang, 2022).
+The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than the model without prejudiced firms.
 This suggests more volatile procyclical wage and wealth responses for black workers than for white workers in the presence of $p$ firms.
-Lastly, we don’t observe a discerning difference in the consumption responses with and without $p$ firms.
+Moreover, consumption among both black and white workers responds more strongly in the benchmark model, leading to a more volatile consumption reaction following the expansionary TFP shock.
 5.3.2 Stochastic simulation
-We simulate the stochastic equilibrium of the economy over 10,000 periods under aggregate TFP shocks, comparing outcomes with and without $p$ firms.
-While impulse responses to an MIT shock capture the transition path following a one-time disturbance, the stochastic equilibrium focuses on the economy’s short-term responses to continuous aggregate shocks.
+We simulate the stochastic equilibrium of the economy over 10,000 periods under aggregate TFP shocks, and compare the outcomes with and without $p$ firms.
+While impulse responses capture the transition path following a one-time shock, stochastic equilibrium focuses on the economy’s short-term responses to a series of aggregate shocks.
 The results are summarized in Table 8.
-In a counterfactual equilibrium without $p$ firms, the racial unemployment rate gap closes, eliminating its cyclical characteristics.
-Likewise, the disparities in the cyclicality and volatility of average wages between racial groups are significantly diminished.
-According to Table 8, the benchmark model reveals a pronounced countercyclicality (-0.6731) and volatility (0.3009) for the black-white unemployment rate gap, both of which disappear when $p$ firms are not included.
-For average wages, the cyclicality for black workers lowers from 1.0749 in the benchmark to 1.0131 without $p$ firms, whereas it increases slightly for white workers from 0.9984 to 1.0115.
-Conversely, the volatility for black workers drops from 1.0755 to 1.0139, while for white workers, it rises from 0.9985 to 1.0116.
-This convergence indicates that hiring discrimination significantly contributes to the cyclical disparities in the labor market between black and white workers.
-These results are consistent with Cajner et al.
-(2017), Ganong et al.
-(2020), and Ragusett (2022), which argue that demographic differences alone do not sufficiently explain the racial disparities observed throughout the business cycle.
-Table 8: Impact of hiring discrimination on business cycle racial gaps
-Cyclicality Volatility
+The stochastic simulation results further support the insights from the impulse response analysis.
+We find that the presence of $p$ firms increases unemployment rate volatility for both black and white workers, but the effect is more pronounced for black workers.
+In the counterfactual economy without $p$ firms, both groups have an unemployment rate relative volatility of 0.18.
+With $p$ firms in the benchmark economy, the volatility for black workers rises to 0.26, compared to 0.22 for white workers.
+Similarly, the volatility of the black-white unemployment rate gap increases from 0.17 to 0.30, reflecting the disproportionate impact on black workers.
+The wages of both black and white workers have dynamics similar to their unemployment rates.
+In particular, the presence of $p$ firms in the benchmark increases both the cyclicality and volatility of wages compared to the counterfactual without $p$ firms.
+However, this effect is more pronounced for black workers than for white workers.
+In our comparison of the benchmark and counterfactual economies, the inclusion of $p$ firms raises the cyclicality and volatility of black wages from 1.01 to 1.08, while white wages increase from 0.98 to 1.
+Overall, hiring discrimination not only heightens wage fluctuations but also widens the gap between black and white workers.
+These results support the findings of Cajner et al.
+(2017), Boulware and Kuttner (2019, 2024), Ganong et al.
+(2020), and Ragusett (2022), who argue that demographic differences alone cannot account for the racial disparities observed over the business cycle.
+Table 8: Impact of hiring discrimination over business cycle
+Volatility Cyclicality
 Benchmark No $p$ firms Benchmark No $p$ firms
 Unemployment rate
-Black-White gap -0.6731 0.0000 0.3009 0.0000
+Black 0.26 0.18 -0.69 -0.70
+White 0.22 0.18 -0.69 -0.70
+Black-White gap 0.30 0.17 -0.67 -0.71
 Average wage
-Black 1.0749 1.0131 1.0755 1.0139
-White 0.9984 1.0115 0.9985 1.0116
-Wealth ratio
-Wealth ratio 0.2028 0.2054 0.1611 0.1466
-Consumption ratio
-Consumption ratio 0.9717 0.9695 0.1861 0.1659
-Note: The top panel compares the black-white unemployment rates, wages, consumption, and wealth gaps over the business cycles between the benchmark model and the model without $p$ firms.
-We further examine the cyclical properties of the average black-to-white wealth ratio and consumption ratio.
-In the benchmark model, the cyclicality of the average wealth ratio is 0.2028, which changes slightly to 0.2054 without $p$ firms, while its volatility decreases from 0.1611 to 0.1466.
-For the average consumption ratio, the cyclicality decreases marginally from 0.9717 to 0.9695, and the volatility drops more significantly from 0.1861 to 0.1659.
-These results show that, aside from the cyclicality of the wealth ratio, the cyclical properties of racial gaps across all other dimensions are notably reduced when discriminatory hiring is eliminated.
-In summary, $p$ firms drive racial disparities in the labor market over the business cycle.
-The black-white unemployment rate gap, countercyclical in the benchmark, vanishes without $p$ firms.
-Wage cyclicality and volatility for black workers also decline, converging toward those of white workers.
-While the wealth ratio’s cyclicality rises slightly, its volatility and that of the consumption ratio decrease.
-These findings indicate that hiring discrimination is a key factor shaping racial differences in the cyclical properties of the labor market and economic outcomes.
-5.4 Welfare disparities in business cycles
-Lastly, we examine the business cycle welfare implications of removing hiring discrimination.
-We simulate the economy for 10,000 periods and compute the long-run average welfare for each individual type, following Cho and Ma (2023).
-The welfare measure is calculated as average consumption equivalence described in Section 4.3 over the simulation periods.
-We present the results in Table 9.
-Table 9: Heterogeneous welfare change in business cycles
-Average welfare gain (%) Eliminate $p$ firms
-Overall -0.01
-Black White
-Average 0.02 -0.01
-by productivity
-Low 0.00 -0.01
-Mid 0.02 -0.01
-High 0.02 -0.02
-by wealth
-Low 20% 0.01 -0.01
-40-60% 0.02 -0.02
-Top 20% 0.03 -0.02
-Note: This table compares the average consumption equivalence change from the benchmark model to the one without $p$ firms in business cycles.
-We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type.
-Low, mid, and high productivity correspond to the lowest, middle, and highest values of idiosyncratic productivity $s$.
-The wealth quintiles are based on benchmark steady-state wealth distribution.
-The economy experiences an average welfare decrease of 0.01% after removing prejudiced firms.
-This is much less than the welfare change in the steady state in Table 5, primarily because workers are risk averse.
-All workers experience a decrease in unemployment rate volatility, while white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions.
-On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%.
-Among white workers, those with high productivity and those in the top 20% of the wealth distribution lose the most (0.02%).
-In contrast, black workers in the top wealth quintile benefit the most, with an average welfare gain of 0.03%.
+Black 1.08 1.01 1.08 1.01
+White 1.00 0.98 1.00 0.98
+Average wealth
+Black 0.39 0.38 0.12 0.12
+White 0.24 0.24 0.07 0.07
+Wealth ratio 0.16 0.14 0.21 0.22
+Average consumption
+Black 0.57 0.52 0.95 0.94
+White 0.39 0.36 0.93 0.92
+Consumption ratio 0.18 0.17 0.97 0.97
+Note: This table compares the responses of black-white unemployment rates, wages, consumption, and wealth gaps over the business cycle between the benchmark model and the model without $p$ firms.
+We further examine the cyclical properties of wealth and consumption for black and white workers.
+Consistent with our earlier findings on labor market outcomes, the presence of $p$ firms increases the volatility of both the wealth and consumption ratios.
+Specifically, with $p$ firms, the volatility of the wealth ratio rises from 0.14 to 0.16, and that of the consumption ratio increases from 0.17 to 0.18.
+These results indicate that, beyond labor market variables, the volatility of racial gaps in other dimensions is also notably reduced when discriminatory hiring is eliminated.
+In summary, our analysis shows that hiring discrimination amplifies racial disparities over the business cycle.
+The presence of $p$ firms increases the volatility of outcomes for black workers relative to white workers across multiple dimensions, including labor market variables, wealth, and consumption.
+These findings suggest that discriminatory hiring plays a key role in driving the cyclical differences in economic outcomes between racial groups.
 6 Conclusion
-This study examines the effects of racial hiring discrimination within a frictional labor market on employment, wage disparities, and wealth accumulation between black and white workers in the U.S.
-We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation.
+This study examines the effects of discrimination in the hiring of black workers on employment, wage disparities, and wealth accumulation in the US.
+We develop a search-and-matching model featuring firms with and without racial prejudices.
 Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in the steady state and business cycles.
-Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum.
-Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome within frictional markets.
-Eliminating these discriminatory factors results in a decrease in overall welfare.
-Our findings shed light on the enduring nature of Black-White racial disparities in the U.S., offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
+In addition, discriminatory hiring places black workers disproportionately at the lower end of the wealth spectrum.
+Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome in frictional labor markets.
+Further, eliminating discriminatory hiring results in an increase in overall economic welfare.
+Our findings shed light on the enduring nature of black-white racial disparities in the US, offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
 Appendix A Stationary equilibrium
@@
 Appendix B Comparing the impact of bargaining power and extreme wealth shock
-Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks.
-Equalizing black workers’ bargaining power with that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes.
-This is because non-prejudiced firms retain less profit, hence post fewer vacancies.
-On the net, it also reduces white workers’ job-finding rate and average wage rate.
-As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
+Our model separately calibrates racial differences in bargaining power and the likelihood of extreme wealth destruction - features we refer to as non-market racial disparities.
+We conduct additional counterfactual analyses to assess the impact of these non-market factors relative to hiring discrimination.
+Table B.1 compares the benchmark model to versions that eliminate non-market disparities by equalizing bargaining power and extreme wealth shocks across racial groups.
+When we set the bargaining power of black workers equal to that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$), their bargained wages rise directly.
+However, the job-finding rate for black workers falls.
+This outcome occurs because non-prejudiced firms retain lower profits and, therefore, post fewer vacancies, which also reduces the job-finding rate and average wage for white workers.
+Consequently, the average wage gap shrinks to 20% (a wage ratio of 80%), accounting for roughly 20% of the observed racial wage gap.
 Table B.1: Aggregate impact of racial disparities from non-market factors
@@
 job separation rate - white 0.04 0.04 0.04
-job finding rate - black 0.50 0.43 0.48
+job finding rate - black 0.49 0.44 0.48
 job finding rate - white 0.66 0.64 0.66
@@
 unemp rate - white 0.05 0.05 0.05
-mean wage - black 1.60 1.71 1.66
-mean wage - white 2.13 2.12 2.15
-mean wage ratio 0.75 0.81 0.77
-mean wealth ratio 0.28 0.29 0.77
+mean wage - black 1.59 1.70 1.65
+mean wage - white 2.12 2.11 2.13
+mean wage ratio 0.75 0.80 0.77
+mean wealth ratio 0.29 0.30 0.77
 median wealth ratio 0.33 0.33 0.88
 Firms
-$p$ firm profit 0.04 0.04 0.04
-np firm profit 0.03 0.02 0.02
+$p$ firm profit 0.04 0.05 0.04
+np firm profit 0.02 0.02 0.02
 p firm vacancy 0.02 0.02 0.02
-np firm vacancy 0.05 0.04 0.05
-p firm employment 0.54 0.58 0.56
-np firm employment 0.40 0.35 0.38
+np firm vacancy 0.05 0.04 0.04
+p firm employment 0.57 0.61 0.58
+np firm employment 0.37 0.33 0.35
 Labor Market
-$p$ market tightness 0.40 0.46 0.43
-$np$ market tightness 0.73 0.60 0.69
+$p$ market tightness 0.39 0.44 0.42
+$np$ market tightness 0.71 0.60 0.67
 Aggregate Outcomes
-Y 3.06 3.05 3.08
-K/Y 10.26 10.25 10.45
-average wage 2.04 2.05 2.06
-unemp rate 0.07 0.07 0.07
+Y 3.07 3.06 3.08
+K/Y 10.26 10.25 10.40
+average wage 2.05 2.05 2.06
+unemp rate 0.06 0.06 0.06
 Note: This table compares the benchmark steady state to a model with equal bargaining power $\xi_{bl}=\xi_{wh}$ and a model with equal extreme wealth shock $\epsilon_{bl}=\epsilon_{wh}$.
-In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers.
+In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same extreme wealth shock as white workers, an external condition governing the accumulating wealth.
 The effect resembles assigning higher bargaining power to black workers.
-This is due to the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012).
+This is due to the importance of personal wealth in helping workers self-insurance against uncertain adverse outcomes (Nakajima, 2012).
 With a lower probability of losing their wealth, black workers can accumulate more personal wealth.
 Higher personal wealth gives black workers a higher reservation value when bargaining with firms.
-Effectively, black workers can negotiate for better wage outcomes (1.66 compared to 1.60 in the benchmark model).
-Similar to directly assigning higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent.
-The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output.
-It spills over to an increase in the average wage rate for all.
-Though it has a much lesser impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap.
-Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).
-Appendix C Impulse responses
+Effectively, black workers can negotiate for better wage outcomes (1.65 compared to 1.59 in the benchmark model).
+Similar to having higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent.
+The lower extreme wealth destruction rate also translates to higher aggregate capital accumulation and increased aggregate output.
+It spills over to an increase in the average wage rate for all, raising the average black-white wage ratio to 77%, thus explaining around 8% of the racial wage gap.
+Since it directly raises the wealth position of black workers, the mean and median black-white wealth ratios increase drastically (77% and 88% from the benchmark levels of 29% and 33%).
+Appendix C Welfare calculation
+We follow Krusell et al.
+(2010) in measuring the welfare effects of discriminatory hiring on black and white workers.
+We calculate the consumption equivalence in a counterfactual scenario without $p$ firms.
+Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$.
+In an alternative economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state.
+We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$.
+Under $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$.
+We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change.
+We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers from the states associated with $p$ firms.
+Aggregating using the benchmark distribution overstates the welfare change for white workers without accounting for this distribution shift.
+Appendix D Measuring the business cycle properties for wage
+Aggregate data can be misleading when estimating wage cyclicality because shifts in worker composition over the business cycle tend to overweight lower-skilled (and lower-wage) workers during expansions, resulting in countercyclical bias.
+To overcome this issue, we use individual-level PSID data (1996-2014), which allows us to hold the composition of workers fixed over time.
+Following the methodology of Bils (1985), Solon et al.
+(1994), and Devereux (2001), we estimate wage cyclicality by regressing the change in individual log wages on the change in log GDP while controlling for key demographic characteristics.
+Our specification is as follows:
+$\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it},$ (D.1)
+where $Y_{t}$ is real GDP.
+The demographic variables $x_{it}$ include gender, education, labor market experience, and marital status.
+We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$.
+In this regression, the coefficient $\beta$ captures the cyclicality of wages.
+We then compute wage volatility as the standard deviation of wage growth after controlling for the demographics.
+The year fixed effects from the regression reflect aggregate wage growth and provide a measure of this volatility.
+This approach ensures that our estimates of wage cyclicality and volatility reflect true economic fluctuations rather than changes in workforce composition.
+Appendix E Impulse responses
 We follow Boppart et al.
@@
 This method also requires the business cycle dynamics to be well approximated as a linear system.
-Figure C.1 presents the impulse response functions for the benchmark model to a 1% positive and negative TFP shock at date 0.
+Figure E.1 presents the impulse response functions for the benchmark model to a 1% positive and a 1% negative TFP shock at date 0.
 The shock gradually returns to a steady state with a persistence of 0.95.
-Our model produces symmetric impulse responses across all variables.
-Figure C.1: Comparing impulse responses
+Our model produces symmetric impulse responses for positive and negative TFP shocks for all variables.
+Figure E.1: Comparing impulse responses
 Note: This figure plots the impulse response functions of the benchmark economy to a one percent increase and a one percent decrease in aggregate TFP at date zero.